### v1

# A Comprehensive Review on Harnessing Large Language Models to Overcome Recommender System Challenges

Abstract.
Recommender systems have traditionally followed modular architectures comprising candidate generation, multi-stage ranking, and re-ranking, each trained separately with supervised objectives and hand-engineered features. While effective in many domains, such systems face persistent challenges including sparse and noisy interaction data, cold-start problems, limited personalization depth, and inadequate semantic understanding of user and item content. The recent emergence of Large Language Models (LLMs) offers a new paradigm for addressing these limitations through unified, language-native mechanisms that can generalize across tasks, domains, and modalities. In this paper, we present a comprehensive technical survey of how LLMs can be leveraged to tackle key challenges in modern recommender systems. We examine the use of LLMs for prompt-driven candidate retrieval, language-native ranking, retrieval-augmented generation (RAG), and conversational recommendation, illustrating how these approaches enhance personalization, semantic alignment, and interpretability without requiring extensive task-specific supervision. LLMs further enable zero- and few-shot reasoning, allowing systems to operate effectively in cold-start and long-tail scenarios by leveraging external knowledge and contextual cues. We categorize these emerging LLM-driven architectures and analyze their effectiveness in mitigating core bottlenecks of conventional pipelines. In doing so, we provide a structured framework for understanding the design space of LLM-enhanced recommenders, and outline the trade-offs between accuracy, scalability, and real-time performance. Our goal is to demonstrate that LLMs are not merely auxiliary components but foundational enablers for building more adaptive, semantically rich, and user-centric recommender systems
This work does not relate to position at Linkedin, Meta, Amazon.
1. Introduction
Recommender systems have become essential across a broad spectrum of digital applications, including content streaming, e-commerce, education, recruiting, and social media platforms (Gomez-Uribe et al., 2022; Cheng et al., 2023b; Agarwal et al., 2019). From personalized playlists on Spotify to tailored learning paths in MOOCs and targeted ads on LinkedIn, recommender systems play a pivotal role in shaping user experience and driving engagement. Despite their widespread adoption and continuous improvements, these systems face enduring challenges such as data sparsity, cold-start problems, dynamic user interests, and explainability (Panigrahi et al., 2023). As modern pipelines grow in complexity—with multi-stage architectures, large-scale retrieval, and diverse modalities—the demands for scalability, transparency, and adaptability have intensified. Recommender systems have traditionally followed modular architectures comprising candidate generation, multi-stage ranking, and re-ranking—each trained separately using supervised objectives and hand-engineered features  (Chen et al., 2023c). While effective in many domains, such systems face persistent challenges, including sparse and noisy interaction data, cold-start problems, limited personalization depth, and inadequate semantic understanding of user and item content.
The recent emergence of Large Language Models (LLMs) offers a new paradigm for addressing these limitations through unified, language-native mechanisms that can generalize across tasks, domains, and modalities. In this paper, we present a comprehensive technical survey of how LLMs can be leveraged to tackle key challenges in modern recommender systems. We examine the use of LLMs for prompt-driven candidate retrieval, language-native ranking, retrieval-augmented generation (RAG), and conversational recommendation—illustrating how these approaches enhance personalization, semantic alignment, and interpretability without requiring extensive task-specific supervision. LLMs also enable zero- and few-shot reasoning, allowing systems to operate effectively in cold-start and long-tail scenarios by leveraging external knowledge and contextual cues. We categorize these emerging LLM-driven architectures and analyze their effectiveness in mitigating core bottlenecks of conventional pipelines. In doing so, we provide a structured framework for understanding the design space of LLM-enhanced recommender systems, and outline trade-offs between accuracy, scalability, and real-time performance.
Our work provides a foundation for researchers and engineers to reimagine recommender systems in the era of large language modeling, and highlights promising directions for future innovation.
2. Evolution of Architectures in Recommender Systems
The architectural landscape of recommender systems has undergone a remarkable transformation over the past two decades. Beginning with simple heuristic methods and collaborative filtering techniques, the field has progressively embraced more sophisticated paradigms powered by deep learning, graph neural networks, and, more recently, LLMs. This evolution reflects the growing demand for personalization, scalability, and contextual awareness in real-world applications. This section presents a chronological overview of key architectural advancements in recommender systems, highlighting the motivations, technical foundations, and practical implications of each generation. By tracing this progression, we aim to shed light on how the field has adapted to increasingly complex data modalities, user behaviors, and computational constraints.
2.1. Traditional Approaches
Heuristic-Based: Heuristic-based recommender systems represent some of the earliest strategies for generating item suggestions. These approaches rely on deterministic rules and statistical aggregations derived from historical data, rather than modeling latent user-item interactions. They are often used in large-scale systems for bootstrapping recommendations or serving as baseline models (Schafer et al., 2001; Bobadilla et al., 2018). A common heuristic is popularity-based recommendation, which ranks items based on global interaction frequency. For instance, the recommendation score $s_{i}$ for an item $i$ may be computed as:
$s_{i}=\text{count}(i)=\sum_{u\in\mathcal{U}}\mathbb{I}(i\in I_{u})$
where $\mathbb{I}(\cdot)$ is the indicator function, $I_{u}$ is the set of items interacted with by user $u$, and $\mathcal{U}$ is the set of all users. This formulation favors frequently interacted items, such as most-viewed videos or best-selling products.
Another widely used rule is co-occurrence-based recommendation, which suggests items based on observed pairwise co-interactions. For example, an associative rule like “Users who bought or viewed item $i$ also bought or viewed item $j$” is operationalized using:
$\text{sim}(i,j)=\frac{\text{count}(i,j)}{\sqrt{\text{count}(i)\cdot\text{count}(j)}}$
where $\text{count}(i,j)$ is the number of users who interacted with both items $i$ and $j$, serving as a symmetric co-occurrence similarity metric (Sarwar et al., 2001). Despite their simplicity and scalability, heuristic methods are inherently non-personalized—they assume user homogeneity and do not adapt to individual preferences. As such, they suffer from popularity bias and lack context-awareness, making them less suitable for tasks requiring fine-grained personalization (Adomavicius and Tuzhilin, 2005). Nonetheless, their low computational cost and interpretability make them useful in cold-start scenarios where user history is sparse, in bootstrapping candidate sets in multi-stage recommender pipelines, and in serving fallback recommendations when more advanced model outputs are unreliable. In many production systems, heuristic scores are combined with learned models in later ranking stages or incorporated into ensemble frameworks for robustness.
Figure 1. (i) Content-Based Recommendation: Recommends movies based on content features such as genre and cast similarity with previously watched movies. (ii) User-Based Collaborative Filtering: Recommends movies by identifying users with similar tastes and suggesting items liked by similar users. (Mallick, 2023)
Collaborative Filtering: Collaborative Filtering (CF)  (Zhang et al., 2014) is a core methodology in recommender systems that predicts a user’s interest in an item by leveraging patterns of user-item interactions. The central assumption behind CF is the homophily principle: users who have shown similar interaction patterns in the past are likely to share future preferences. Unlike content-based methods, CF requires no side information or item/user metadata—it exclusively relies on interaction history, such as ratings, clicks, or purchases.
Formally, given a sparse interaction matrix $\mathbf{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$ where $\mathcal{U}$ is the set of users and $\mathcal{I}$ is the set of items, the goal is to estimate the unobserved ratings or affinities $\hat{r}_{u,i}$ for a user-item pair $(u,i)$.
There are two major paradigms of CF. User-based CF computes the similarity between the target user $u$ and other users $v\in\mathcal{U}$, and then aggregates the preferences of the $k$ most similar users $\mathcal{N}_{k}(u)$ to predict a score using the formula:
$\hat{r}_{u,i}=\frac{\sum_{v\in\mathcal{N}_{k}(u)}\text{sim}(u,v)\cdot r_{v,i}}{\sum_{v\in\mathcal{N}_{k}(u)}|\text{sim}(u,v)|}$
where $\text{sim}(u,v)$ may be computed using cosine similarity, Pearson correlation, or Jaccard similarity over the vectors $\mathbf{r}_{u}$ and $\mathbf{r}_{v}$.
In contrast, item-based CF finds items similar to those the user has interacted with, using item-item similarities. The estimated score is:
$\hat{r}_{u,i}=\frac{\sum_{j\in\mathcal{I}_{u}}\text{sim}(i,j)\cdot r_{u,j}}{\sum_{j\in\mathcal{I}_{u}}|\text{sim}(i,j)|}$
where $\mathcal{I}_{u}$ is the set of items interacted with by user $u$, and similarity can again be cosine or adjusted cosine similarity.
Common similarity functions include cosine similarity:
$\text{sim}_{\text{cosine}}(i,j)=\frac{\mathbf{r}_{*,i}\cdot\mathbf{r}_{*,j}}{\|\mathbf{r}_{*,i}\|\cdot\|\mathbf{r}_{*,j}\|}$
and Pearson correlation:
$\text{sim}_{\text{Pearson}}(i,j)=\frac{\sum_{u}(r_{u,i}-\bar{r}_{i})(r_{u,j}-\bar{r}_{j})}{\sqrt{\sum_{u}(r_{u,i}-\bar{r}_{i})^{2}}\sqrt{\sum_{u}(r_{u,j}-\bar{r}_{j})^{2}}}$
Despite its simplicity and interpretability, CF faces several challenges. Chief among them are data sparsity and the cold-start problem—both user and item-level—which hinder similarity computation. Real-time similarity computation can also become computationally expensive as the number of users or items grows. To mitigate these limitations, CF is often combined with content-based models in hybrid architectures.
Furthermore, recent advances have proposed learning-based CF methods such as matrix factorization and neural embeddings. For instance, LightGCN (He et al., 2020) simplifies graph-based collaborative filtering by removing feature transformations and non-linearities, using only weighted summation of neighborhood embeddings:
$\mathbf{e}_{u}^{(k)}=\sum_{i\in\mathcal{N}(u)}\frac{1}{\sqrt{|\mathcal{N}(u)||\mathcal{N}(i)|}}\cdot\mathbf{e}_{i}^{(k-1)}$
This iterative propagation enhances higher-order connectivity, improving recommendation accuracy while maintaining linear complexity. Collaborative Filtering remains a fundamental building block in many industrial pipelines and serves as a baseline for benchmarking modern deep learning recommender systems (Li et al., 2023d; Rendle et al., 2020).
Matrix Factorization: Matrix Factorization (MF)  (Zhang, 2022) is a class of latent factor models that projects users and items into a shared low-dimensional vector space to model user-item interactions. Given a user–item rating matrix $R\in\mathbb{R}^{m\times n}$, where $m$ is the number of users and $n$ is the number of items, the goal is to approximate $R$ by the product of two low-rank matrices:
$R\approx\hat{R}=UV^{\top},$
where $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$ are the user and item latent factor matrices, and $k\ll\min(m,n)$ is the number of latent dimensions.
The optimization objective is typically to minimize the squared error over observed entries:
$\min_{U,V}\sum_{(u,i)\in\mathcal{K}}(r_{u,i}-\mathbf{u}_{u}^{\top}\mathbf{v}_{i})^{2}+\lambda\left(\|\mathbf{u}_{u}\|^{2}+\|\mathbf{v}_{i}\|^{2}\right),$
where $\mathcal{K}$ is the set of observed ratings, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent vectors for user $u$ and item $i$, respectively, and $\lambda$ is a regularization hyperparameter to prevent overfitting.
Several variations of matrix factorization have been proposed to enhance recommendation accuracy and scalability. Singular Value Decomposition (SVD)  (Klema and Laub, 1980) applies truncated decomposition to the user-item rating matrix, projecting it into lower-dimensional latent factors; however, it often requires prior imputation of missing entries, which can introduce bias. Alternating Least Squares (ALS) offers a more scalable alternative by optimizing the matrix factorization objective through iterative updates: fixing one of the user or item matrices while solving a regularized least squares problem for the other (Zhou et al., 2008). This alternating procedure makes ALS particularly well-suited for large-scale and distributed environments, such as those in industrial settings. An extension to these approaches, SVD++, incorporates implicit feedback—such as clicks or views—alongside explicit ratings to refine user representations. The predicted rating in SVD++ is given by:
$\hat{r}_{u,i}=\mu+b_{u}+b_{i}+\left(\mathbf{u}_{u}+\sum_{j\in\mathcal{I}_{u}}\frac{y_{j}}{\sqrt{|\mathcal{I}_{u}|}}\right)^{\top}\mathbf{v}_{i},$
where $\mu$ is the global average rating, $b_{u}$ and $b_{i}$ are user and item-specific biases, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent factors for user $u$ and item $i$, respectively, and $y_{j}$ represents the embedding of item $j$ contributing to user $u$’s implicit feedback profile.
Matrix factorization has become the backbone of many collaborative filtering systems due to its balance between expressiveness and scalability. Its popularity surged during the Netflix Prize competition, and it continues to serve as a robust baseline for collaborative filtering tasks (Koren et al., 2009).
2.2. Content-Based and Hybrid Systems
Content-Based Filtering: Content-based filtering (CBF)  (Wang et al., 2024d) recommends items to a user by analyzing the similarity between the items’ attributes and the user’s historical preferences (Lops et al., 2011). Unlike collaborative methods, CBF operates independently of other users and relies solely on item metadata—such as genre, tags, textual descriptions, or visual features.
A user profile $\mathbf{u}$ is typically constructed by aggregating the feature vectors of previously liked or interacted items. For a given item $i$ with a feature vector $\mathbf{x}_{i}$, the predicted preference score is computed using a similarity function:
$\hat{r}_{u,i}=\text{sim}(\mathbf{u},\mathbf{x}_{i})$
Common choices for sim include cosine similarity:
$\text{sim}_{\text{cosine}}(\mathbf{u},\mathbf{x}_{i})=\frac{\mathbf{u}^{\top}\mathbf{x}_{i}}{\|\mathbf{u}\|\cdot\|\mathbf{x}_{i}\|}$
In text-based domains, TF-IDF representations or pre-trained language model embeddings are often used to define $\mathbf{x}_{i}$, while in visual domains, image embeddings from CNNs or vision transformers are leveraged (Mouhiha et al., 2024).
While CBF models are robust to cold-start users (as long as item features are known), they struggle with over-specialization and limited discovery of novel content.
Hybrid Models: Hybrid recommender systems combine content-based and collaborative filtering paradigms to overcome the limitations of either approach in isolation (Burke, 2002). The fusion can occur at various levels—model, feature, or score—and often improves robustness, especially in sparse or cold-start scenarios.
LightFM (Kula, 2015) incorporates both collaborative signals (user-item interactions) and content features (metadata) into a unified latent factor model. Its objective is to learn user and item embeddings from both sources:
$\hat{r}_{u,i}=\mathbf{u}_{u}^{\top}\mathbf{v}_{i}+\mathbf{u}_{u}^{\top}\mathbf{f}_{i}+\mathbf{f}_{u}^{\top}\mathbf{v}_{i},$
where $\mathbf{f}_{u}$ and $\mathbf{f}_{i}$ are feature representations of user and item metadata respectively.
DeepFM (Guo et al., 2017) further extends this by combining a Factorization Machine (FM) layer for low-order feature interactions with a deep neural network for high-order interactions. while the deep component captures non-linear patterns across raw features via multi-layer perceptrons. These hybrid architectures provide stronger generalization across heterogeneous data sources and are widely used in large-scale recommendation engines.
2.3. Deep Learning Architectures
Neural Collaborative Filtering (NCF): Neural Collaborative Filtering (NCF)  (He et al., 2017a) is a deep learning-based extension of matrix factorization that captures complex, nonlinear user–item interaction patterns. Instead of modeling the dot product between latent user and item vectors, NCF uses a Multi-Layer Perceptron (MLP) to learn a nonlinear interaction function:
$\hat{r}_{u,i}=\phi(\mathbf{u}_{u},\mathbf{v}_{i})=\text{MLP}([\mathbf{u}_{u};\mathbf{v}_{i}]),$
where $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the embedding vectors for user $u$ and item $i$, respectively, and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation. The architecture enables the model to approximate arbitrary functions between user and item representations. The NeuMF model (He et al., 2017b) combines Generalized Matrix Factorization (GMF) and MLP-based modeling via a fusion layer to benefit from both linear and nonlinear signals.
Figure 2. Neural Collaborative Filtering (NCF) architecture: One-hot user and item inputs are embedded into latent vectors, passed through neural layers, and produce a predicted score $\hat{y}_{ui}$ trained against the target $y_{ui}$ (He et al., 2017a)
Wide & Deep Models: The Wide & Deep architecture (Cheng et al., 2016) aims to capture both low- and high-order feature interactions by combining a linear model (”wide”) and a deep neural network (”deep”) within a joint framework:
$\hat{y}=\sigma(\mathbf{w}^{\top}\mathbf{x}+\mathbf{a}^{\top}f_{\text{deep}}(\mathbf{x})),$
where $\mathbf{x}$ is the feature vector, $\mathbf{w}$ are the wide (memorization) weights, and $f_{\text{deep}}(\mathbf{x})$ represents the output of the deep network. This hybrid design enables the model to both remember frequent co-occurrences (e.g., user clicked item A and B) and generalize to unseen feature combinations. It was deployed in production at Google Play for app recommendations, demonstrating substantial improvements in personalization and click-through rates.
2.4. Sequence-Aware Models
As user behavior in many applications is inherently sequential (e.g., streaming, shopping, or reading), capturing temporal dependencies becomes crucial. Sequence-aware recommender systems model the order and dynamics of user-item interactions over time, enabling next-item prediction based on behavioral context.
RNN-Based Models: Recurrent Neural Networks (RNNs)  (Schmidt, 2019), particularly Gated Recurrent Units (GRUs) and Long Short-Term Memory (LSTM) networks, have been used to model user behavior sequences. GRU4Rec (Hidasi et al., 2016) was one of the earliest models to apply RNNs to session-based recommendation, treating the problem as a sequence prediction task.
Let $x_{t}$ be the embedding of the item interacted with at time $t$. The GRU updates its hidden state $h_{t}$ using:
$h_{t}=\text{GRU}(x_{t},h_{t-1}),$
and predicts the next item via:
$\hat{y}_{t+1}=\text{softmax}(Wh_{t}+b),$
where $W$ and $b$ are learnable parameters. GRU4Rec enables learning from both short-term (within-session) and long-term user preferences and has shown superior performance in session-based scenarios compared to traditional CF methods.
Transformer-Based Models: Transformer-based  (Vaswani et al., 2023) models replace the recurrence in RNNs with self-attention mechanisms, allowing better modeling of long-range dependencies and parallelization. SASRec (Kang and McAuley, 2018) applies a unidirectional self-attention encoder to model sequential interactions:
$\mathbf{H}=\text{SelfAttention}(\mathbf{X}+\mathbf{P}),$
where $\mathbf{X}$ is the sequence of item embeddings and $\mathbf{P}$ is the positional encoding. The model predicts the next item using the last hidden state:
$\hat{y}_{t+1}=\text{softmax}(Wh_{t}).$
BERT4Rec (Sun et al., 2019) introduces bidirectional attention with a masked item prediction objective, inspired by the BERT architecture in NLP. This enables the model to learn contextual dependencies from both past and future items in a sequence.
TiSASRec (Li et al., 2020) extends SASRec by incorporating both absolute and relative time intervals into attention weights, enhancing temporal modeling in user behavior:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}+T_{\text{abs}}+T_{\text{rel}}}{\sqrt{d_{k}}}\right)V,$
where $T_{\text{abs}}$ and $T_{\text{rel}}$ encode absolute and relative temporal distances, respectively.
Overall, transformer-based models have surpassed RNNs in many sequence-aware recommendation tasks due to their ability to model long-term dependencies and leverage parallel computation.
2.5. Graph-Based Recommenders
Graph-based recommender systems  (Wang et al., 2021a) represent user–item interactions as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathcal{E})$, where $\mathcal{U}$ is the set of users, $\mathcal{I}$ is the set of items, and edges $\mathcal{E}$ correspond to observed interactions (e.g., clicks, ratings, purchases). This graph structure enables the modeling of high-order connectivity and collaborative signals beyond immediate neighbors, capturing richer semantics than traditional matrix factorization.
Figure 3. Illustration of a heterogeneous graph-based recommender system integrating three components: the Social Relation Graph (user-user links), the User-Item Interaction Graph (explicit and implicit feedback such as clicks and ratings), and the Domain Knowledge Graph (item attributes like OS and color). This structure facilitates richer context-aware and socially-informed recommendations (Kipf and Welling, 2017)
Graph Convolutional Networks (GCNs) have been widely adopted for such modeling  (Kipf and Welling, 2017). A GCN propagates information between nodes through message passing, updating the representation of a node by aggregating features from its neighbors. In a simplified GCN layer:
$\mathbf{h}_{v}^{(k)}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{W}^{(k)}\mathbf{h}_{u}^{(k-1)}\right),$
where $\mathbf{h}_{v}^{(k)}$ is the embedding of node $v$ at the $k^{\text{th}}$ layer, $\mathbf{W}^{(k)}$ is a learnable weight matrix, and $\mathcal{N}(v)$ is the set of neighbors of $v$.
NGCF (Neural Graph Collaborative Filtering) (Ying et al., 2018) builds on this idea by combining collaborative filtering with graph convolution. It explicitly models user–item interactions via non-linear message propagation. At each layer, it computes:
$\mathbf{e}_{v}^{(k)}=\sigma\left(\mathbf{W}_{1}^{(k)}\mathbf{e}_{v}^{(k-1)}+\mathbf{W}_{2}^{(k)}\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}\right),$
where $\mathbf{e}_{v}^{(k)}$ is the embedding of user or item $v$ at layer $k$.
LightGCN (He et al., 2020) simplifies NGCF by removing the nonlinear activation and transformation weights, arguing that they may over-complicate the propagation. LightGCN computes embeddings as:
$\mathbf{e}_{v}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{v}^{(k)},\quad\text{where}\quad\mathbf{e}_{v}^{(k)}=\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}.$
PinSage (Ying et al., 2018) extends this concept for large-scale web-scale recommendation systems (e.g., Pinterest). It combines random walks with graph convolutions and incorporates node features using neighborhood sampling and importance-based aggregation.
Graph-based models are especially effective at capturing high-order collaborative signals and performing inductive reasoning on unseen nodes. They have become foundational in modern recommenders deployed in large-scale platforms due to their flexibility and scalability.
2.6. Large-Scale Retrieval Models
Large-scale retrieval models are designed to efficiently handle recommendation tasks at industrial scale, where the corpus of items may reach millions or billions. These models prioritize fast candidate generation and scalability while maintaining reasonable recommendation quality.
Figure 4. Hybrid ranking architecture combining two-tower and pairwise models. User-specific item representations and shallow embeddings are processed through fusion scoring to generate the final ranked list.  (Yuan et al., 2024).
Two-Tower Models: Two-tower architectures, also known as dual-encoder models  (Yuan et al., 2024), are among the most commonly deployed retrieval methods in production systems. These models learn independent embedding functions for users and items:
$\mathbf{e}_{u}=f_{\theta}(u),\quad\mathbf{e}_{i}=g_{\phi}(i)$
where $f_{\theta}$ and $g_{\phi}$ are neural networks parameterized by $\theta$ and $\phi$, respectively. The final relevance score between user $u$ and item $i$ is computed as the inner product or cosine similarity:
$s(u,i)=\langle\mathbf{e}_{u},\mathbf{e}_{i}\rangle$
This separation allows precomputing item embeddings and indexing them via Approximate Nearest Neighbor (ANN) techniques, enabling sublinear retrieval times.
The YouTube DNN model (Covington et al., 2016) is an influential example of a two-tower retrieval system. It uses multi-layer perceptrons (MLPs) for both user and item towers, trained with sampled softmax to optimize for the target watch history. Another widely used architecture is Facebook’s Deep Learning Recommendation Model (DLRM) (Naumov et al., 2019), which combines dense and sparse features from user-item pairs and leverages embedding lookups and interaction layers to compute the final relevance score.
These models form the first stage in multi-stage ranking pipelines. After candidate generation via ANN search over user-item embeddings, a more expressive model (e.g., gradient boosting, transformers) typically performs re-ranking over the retrieved shortlist.
Due to their efficiency and deployment advantages, two-tower retrieval models are considered foundational in modern industrial recommender systems.
3. Key Components of Recommender System Pipelines
3.1. Candidate Generation
Candidate generation is the first stage in large-scale recommender systems, responsible for filtering a massive corpus of items down to a smaller subset of potentially relevant candidates for a given user. This stage prioritizes recall over precision, ensuring that relevant items are included in the candidate set, even if some irrelevant items are also retrieved. Precision is handled in later stages through more computationally intensive ranking models. The goal is to efficiently retrieve a set of items $\mathcal{C}(u)\subset\mathcal{I}$ for a user $u\in\mathcal{U}$, where $\mathcal{I}$ denotes the entire item corpus. Formally, candidate generation can be defined as retrieving all items $i\in\mathcal{I}$ such that the similarity between the user embedding $f_{u}$ and the item embedding $f_{i}$ exceeds a threshold $\tau$, i.e.,
$\mathcal{C}(u)=\{i\in\mathcal{I}\mid\text{sim}(f_{u},f_{i})\geq\tau\},$
where $\text{sim}(\cdot,\cdot)$ is a similarity function such as dot product or cosine similarity. Several techniques are commonly employed in candidate generation. Approximate Nearest Neighbor (ANN) search allows for efficient similarity-based retrieval using indexing structures like HNSW or product quantization (Malkov and Yashunin, 2020; Jegou et al., 2011). Graph-based methods leverage user-item interaction graphs, using algorithms such as random walks (Perozzi et al., 2014) or personalized PageRank (Haveliwala, 2002) to identify semantically or behaviorally related items. Heuristic filtering approaches use hand-crafted rules based on metadata, popularity, or recent activity to shortlist items quickly. Embedding-based retrieval methods, often built using two-tower or dual encoder models, learn representations for users and items such that similar entities lie close in a shared vector space (Yang et al., 2020). Additionally, session-based models take into account recent user interactions and temporal patterns, often utilizing RNNs (Hidasi et al., 2016) or attention mechanisms (Li et al., 2017) to generate context-aware candidates.
Effective candidate generation requires balancing computational efficiency with the need for broad item coverage. Systems must be optimized for low latency and high throughput while ensuring that the candidate pool includes a diverse and relevant set of items to support personalization downstream. Periodic updates to item indices and embeddings are essential to maintain temporal relevance. Overall, the success of downstream ranking models is closely tied to the quality and coverage of the candidate set produced in this stage.
3.2. Scoring and Ranking
Once the candidate set is generated, the next stage in the recommendation pipeline involves computing a relevance score for each candidate and producing a ranked list. This scoring process is handled by machine learning models that are typically more expressive and computationally expensive than those used in candidate generation, since they operate on a smaller set of items. These models ingest high-dimensional features encompassing user profiles, item attributes, and contextual signals such as time of day, device type, and recent interactions. Common architectures for this stage include gradient-boosted decision trees (GBDT) (Ke et al., 2017), deep neural networks (DNNs) (Covington et al., 2016), and transformer-based models (Sun et al., 2019), which are trained to predict item relevance based on learned representations of user-item pairs.
Formally, for each user $u\in\mathcal{U}$ and candidate item $i\in\mathcal{C}(u)$, the system learns a scoring function $s(u,i)$ defined as:
$s(u,i)=f_{\theta}(\phi_{u},\phi_{i},\psi_{u,i}),$
where $\phi_{u}$ and $\phi_{i}$ are the embedding representations of the user and item, respectively, and $\psi_{u,i}$ denotes additional features capturing user-item interaction context. The function $f_{\theta}$ is typically a feed-forward neural network or transformer-based encoder that outputs a scalar relevance score. The top-ranked items are selected by sorting candidates in descending order of $s(u,i)$.
In multi-task learning settings, the model predicts several outcomes simultaneously—such as click probability, dwell time, and engagement—using a shared encoder and task-specific output heads:
$h=\text{Encoder}_{\theta}(\phi_{u},\phi_{i},\psi_{u,i}),$
$\hat{y}_{\text{click}}=\sigma(W_{\text{click}}h),\quad\hat{y}_{\text{dwell}}=\text{ReLU}(W_{\text{dwell}}h),\quad\hat{y}_{\text{engage}}=\sigma(W_{\text{engage}}h),$
where $h$ is the latent representation, $\sigma$ denotes the sigmoid function, and $W$’s are task-specific projection matrices. The total training loss is computed as a weighted combination of individual task losses:
$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{\text{click}}+\lambda_{2}\cdot\mathcal{L}_{\text{dwell}}+\lambda_{3}\cdot\mathcal{L}_{\text{engage}}.$
This setup enables joint optimization across objectives and improves the model’s ability to learn shared behavioral patterns. Domain-specific features further enhance personalization, such as temporal trends, item freshness, category information, and content embeddings. The scoring and ranking stage is critical for final recommendation quality, as it ultimately determines the ordering and presentation of content to the user.
3.3. Post-Ranking and Calibration
After the initial scoring and ranking phase, post-ranking modules are applied to adjust the ranked list in order to align recommendations with broader system-level goals beyond user-item relevance. These objectives include promoting content diversity, controlling item fatigue, ensuring fairness across item categories or creators, and maintaining freshness in temporal contexts. This stage is critical to shaping the final user experience and achieving long-term business metrics such as content exposure, retention, and platform health.
The input to this module is an already ranked list $\mathcal{R}(u)=\{i_{1},i_{2},\dots,i_{k}\}$ with associated relevance scores $s(u,i_{j})$. Post-ranking adjusts these scores via a transformation function $g(\cdot)$, which considers additional factors such as diversity penalties or freshness boosts:
$s^{\prime}(u,i_{j})=g(s(u,i_{j}),\delta(i_{j}),\gamma(i_{j})),$
where:
  • $s(u,i_{j})$ is the original relevance score,
  • $\delta(i_{j})$ is a diversity-aware adjustment factor (e.g., based on item similarity or content clusters),
  • $\gamma(i_{j})$ is a freshness or recency factor.
One common technique is diversity-aware re-ranking, which penalizes redundant content in the top positions using a marginal gain function. This can be formalized using Determinantal Point Processes (DPP) (Kulesza and Taskar, 2012) or greedy submodular optimization (Ashkan et al., 2015). Similarly, fatigue control mechanisms reduce the score of items already seen or interacted with recently, preventing oversaturation.
In dynamic settings, exploration-exploitation trade-offs are addressed through bandit algorithms (Li et al., 2010) or reinforcement learning (Chen et al., 2019). These methods adjust the final ranking to allow new or uncertain items to be exposed to users with non-zero probability. In a contextual bandit framework, the post-ranking score $s^{\prime}$ may incorporate expected reward:
$s^{\prime}(u,i_{j})=\mathbb{E}[r_{u,i_{j}}\mid x_{u,i_{j}}],$
where $r_{u,i_{j}}$ is the reward (e.g., click, engagement) and $x_{u,i_{j}}$ are the contextual features of the interaction.
Overall, post-ranking ensures that the final list achieves a desirable trade-off between individual relevance and system-level goals. This stage plays a critical role in aligning model outputs with platform constraints and policy requirements while enhancing long-term user satisfaction.
4. Evolving Recommender Systems: Addressing Traditional Challenges with LLM-Based Solutions
Recommender systems deployed in industrial settings face a wide array of challenges, many of which stem from the nature of the data itself. Below, we outline key data-centric issues and describe how leading platforms address them using scalable machine learning techniques, representation learning, and multimodal data fusion.
In this section, we explore how LLMs can be integrated into Recommender system architectures to address six main categories of industrial challenges. We begin with data understanding and cold start issues and then progress through modeling, evaluation, system design, privacy, and organizational concerns. For each area, we discuss not only the strengths of LLM-based approaches but also the practical trade-offs and open research directions they entail.
Table 1. Challenges in Recommender Systems and Corresponding LLM-Based Solutions
Challenges Sub-problem LLM-Based Solutions References
Data-Centric Cold Start Content-Conditioned Generation
Retrieval-Augmented Generation
Zero-Shot Personalization
Representation Bootstrapping
Language-Native Dialogue
Prompt-Based Conditioning
Multimodal Embedding Synthesis
Meta-Learning for Fast Adaptation
Cross-Domain Transfer
(Zhang et al., 2021a; Li et al., 2023a; Zhai, 2024; Xu et al., 2023b; Kang et al., 2023; Dai et al., 2023)
(Huang and Huang, 2024; Gao et al., 2021; Gomez-Uribe and Hunt, 2021; Asai and Hajishirzi, 2020)
(Wei et al., 2022; Purificato et al., 2024; Thomas et al., 2022)
(Tennenholtz et al., 2024; Huber et al., 2025)
(Chan et al., 2024; Wang, 2024; Zheng et al., 2024)
(Ramos et al., 2025; Jiang et al., 2024)
(Ranasinghe et al., 2025; Alayrac et al., 2022; Tian et al., 2025; Yu et al., 2022)
(Wang et al., 2024c; Finn et al., 2017; Lee et al., 2019)
(Vajjala et al., 2024; Yi et al., 2023; Liu et al., 2025b; Vats et al., 2024)
Data Sparsity Text-Driven Generalization
Semantic Matching via Embedding
(Zhang et al., 2021b; Seymour et al., 2023; Hwang and Lee, 2024)
(Wang et al., 2024b; Cui et al., 2024)
Noisy Implicit Feedback Feedback Interpretation with Contextual Prompts (Bodonhelyi et al., 2024; Lian et al., 2020; Lee, 2013; Sun et al., 2017; Zhai, 2024)
Temporal Drift Temporal Adaptation with LLMs
Structural and Hybrid Integration
(Zhao et al., 2025b; Kumar et al., 2025)
(Luo et al., 2024; Han et al., 2025)
Multimodal Data Integration Modal-Aware Weighting
Unified Representation
Cross-Modal Alignment
Multimodal Imputation
Semantic Fusion
(Xu et al., 2023b; Zhai, 2024; Ma et al., 2024)
(Lu et al., 2022, 2023)
(Sapkota et al., 2025; Zhang et al., 2025d)
(Radford et al., 2021; Zhou et al., 2025; Chen et al., 2022)
(Jung and won Lee, 2025; Aghajanyan et al., 2022)
Modeling & Algorithmic Personalization vs. Generalization Instruction-Tuned Generalization
Prompt-Tuned Personalization
Behavioral Diversity
Multitask Prompting
(Liu et al., 2025a)
(Jiang et al., 2024; Li et al., 2022b)
(Wang et al., 2023a)
Scalability of Deep Models LLM-Based Distillation
Prompt-Efficient Inference
Two-Stage Hybrid Pipelines
Sparse Activation Architectures
(Chung et al., 2022; Jiang et al., 2024; Wang et al., 2025a)
(Zhao et al., 2022; Shen et al., 2021; Li et al., 2023a; Liu et al., 2023c)
(Hu et al., 2021; Lester et al., 2021; Zhai, 2024; Ouyang et al., 2022)
(Du et al., 2022; Lin et al., 2021)
Long-Tail Modeling Content-Enriched Generation
Retrieval-Augmented Tail Expansion
Tail-Aware Few-Shot Prompting
Multimodal Tail Representation
(Zhang et al., 2021a; Kang et al., 2023)
(Vasile et al., 2022; Dai et al., 2023; Wu et al., 2024a)
(Chung et al., 2022; Wang et al., 2023a)
(Vasile et al., 2022; Oh et al., 2025)
Evaluation & Experimentation Offline–Online Gap Counterfactual Evaluation
Offline Metric Recalibration
Behavior-Level Satisfaction
Interactive User Simulation
Evaluation Metric Generation
(Gilotte et al., 2018; Saito, 2020; Jeunen and London, 2023)
(Li et al., 2023c; Wu et al., 2024b)
(Zhao et al., 2025a; Lee et al., 2023b)
(Shang et al., 2025; Bemthuis et al., 2024)
(Zhou et al., 2023b; Liu et al., 2023a; Xu et al., 2023a)
Sparse Conversion Labels Proxy Signal Augmentation
Instruction-Tuned Label Imputation
Generative Multi-Task Learning
Counterfactual Label Reasoning
Language-Guided Reweighting
(Dai et al., 2023; Zhai, 2024; Zhang et al., 2021a)
(Wong and Tanaka, 2025)
(Freiberger et al., 2025)
(Joshi et al., 2024)
(Tang et al., 2023)
Balancing Immediate Engagement LLM-Based Proxy Reward Estimation
Counterfactual Dialogue
Preference Drift Detection
Multi-Objective RL with LLM-Guided Reward
(Xue et al., 2022; Park et al., 2019)
(Farshidi et al., 2023; Li et al., 2022a)
(Zheng et al., 2023)
(Zhao et al., 2024; Lee et al., 2023a)
(Christiano et al., 2017; Liu et al., 2023b; Ziegler et al., 2019)
Privacy, Security, & Regulatory User Data Sensitivity Token Attribution and Prompt Fingerprinting (Li et al., 2024b)
Compliance with Laws Streaming Log Replay + Token-Level Attribution (Xie et al., 2023; Park and Chen, 2025)
Differential Privacy & Federated Learning LLMs for Ephemeral Personalization
Gradient-Sanitized Fine-Tuning
Federated Prompting
(McMahan et al., 2017)
(Nguyen and Pan, 2024; Rao and Jin, 2024)
(Li and Roy, 2025; Zhou and Yamamoto, 2025; Wu and Singh, 2025)
Data Access and Governance Data Minimization via LLMs
Instructable Privacy Filters
Synthetic Data Generation
(Bougie and Watanabe, 2025; Liu et al., 2021)
(Chen et al., 2025a; Dong et al., 2024b)
(Xie et al., 2024; Lin et al., 2024)
4.1. Data-Centric Challenges
Recommender systems deployed in industrial environments face a host of challenges rooted not just in algorithmic design, but in the quality, availability, and dynamics of the underlying data. The data-centric issues significantly influence model generalization, fairness, personalization, and system robustness. Below, we dissect key categories of data-centric challenges and discuss how they impact large-scale deployment and learning dynamics in recommender pipelines.
4.1.1. Cold Start Problem
The cold start problem refers to the difficulty of generating reliable recommendations for new users or new items due to the absence of interaction history. This issue is particularly acute in collaborative filtering (CF) systems, where latent representations are learned from observed user–item interactions.
In matrix factorization frameworks, predicted preferences are typically computed as:
$\hat{r}_{ui}=\mathbf{p}_{u}^{\top}\mathbf{q}_{i}$
where $\mathbf{p}_{u},\mathbf{q}_{i}\in\mathbb{R}^{d}$ are the latent embeddings of user $u$ and item $i$. These embeddings are optimized by minimizing a loss over the set of observed interactions $\mathcal{O}$:
$\min_{\{\mathbf{p}_{u}\},\{\mathbf{q}_{i}\}}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}(\hat{r}_{ui},r_{ui})+\lambda\left(\|\mathbf{p}_{u}\|^{2}+\|\mathbf{q}_{i}\|^{2}\right)$
In cold start settings, when $u^{\prime}\notin\mathcal{O}_{u}$ or $i^{\prime}\notin\mathcal{O}_{i}$, the corresponding embeddings $\mathbf{p}_{u^{\prime}}$ or $\mathbf{q}_{i^{\prime}}$ are poorly initialized or undefined, leading to degraded or invalid predictions:
$\hat{r}_{u^{\prime}i}=\mathbf{p}_{u^{\prime}}^{\top}\mathbf{q}_{i}\quad\text{or}\quad\hat{r}_{ui^{\prime}}=\mathbf{p}_{u}^{\top}\mathbf{q}_{i^{\prime}}$
This challenge is exacerbated in large-scale, dynamic environments where thousands of new users and items enter the system daily (Quadrana et al., 2017). Even more complex architectures such as two-tower models (Covington et al., 2016) or deep sequential recommenders (Hidasi et al., 2016) rely on embedding representations learned from prior interactions, and thus remain vulnerable to cold start scenarios. The sparsity of user–item signals and the lack of initialization paths for unseen entities make cold start a persistent bottleneck in the design of effective recommender systems (Schein et al., 2002; Park and Tuzhilin, 2006). LLMs offer a paradigm shift in mitigating the cold start problem by leveraging their ability to process rich textual and contextual information without requiring historical interaction data (Dai et al., 2023; Wang et al., 2024b). Unlike traditional collaborative filtering methods that depend on dense user–item matrices, LLMs can infer relevance through semantic reasoning, prompt-based generation, and cross-modal understanding. Let’s examine how LLMs can solve the cold start problem in recommender systems.
Content-Conditioned Generation – LLMs pretrained on web-scale corpora possess rich world knowledge and linguistic priors, enabling them to generate item recommendations solely from textual descriptions. This paradigm is particularly advantageous in cold-start and low-resource scenarios where interaction histories are sparse or nonexistent. Traditional recommender systems often struggle in these settings, as they depend heavily on collaborative signals (e.g., clicks, ratings, watch time). In contrast, LLMs can interpret and reason over content attributes such as titles, descriptions, reviews, genres, and even product specifications.
For instance, Generative Recommenders (Zhang et al., 2021a; Li et al., 2023a) employ prompt-based methods where user queries or profiles are combined with product metadata to produce ranked outputs or candidate lists. In these setups, the generation process bypasses the need for learned user-item embeddings and instead leverages natural language priors encoded in the LLM:
$\hat{r}_{ui}=\text{LLM}(\texttt{``User profile: ''}+x_{u}+\texttt{`` Item: ''}+x_{i})$
Such models can be further enhanced with instruction prompts, prefix tuning, or adapters to align generation with domain-specific preferences and intents. Recent work demonstrates that even simple prompts can elicit meaningful ranking signals, especially when the model is instructed to output preference scores, reasoning traces, or direct answers.
This approach has been adopted in real-world pipelines, including JD.com’s product recommendation and Pinterest’s item-to-item generation (Zhai, 2024; Xu et al., 2023b), where structured item metadata and textual attributes are encoded as input contexts. Furthermore, systems like GPTRec and PromptRec have shown how generative LLMs can synthesize interaction labels (e.g., likelihood of liking or purchasing an item) from textual descriptions, enabling bootstrap datasets for pretraining, self-training, or teacher-student distillation in downstream rankers.
By conditioning generation on content alone, these models enable rapid personalization without requiring prior interactions, thereby offering a scalable solution to cold-start recommendation, catalog expansion, long-tail item surfacing, and new-user onboarding. As instruction tuning and few-shot prompting mature, we expect content-conditioned generation to play a central role in building adaptable, zero-shot capable recommender agents that generalize across domains with minimal supervision.
Retrieval-Augmented Generation (RAG) – RAG-based methods enhance LLMs with a retrieval module that selects relevant items or documents based on the input query. In recommender systems, this architecture decouples memorization from inference, allowing the model to attend to similar items, behavioral patterns, or metadata even when the target entity is entirely new or sparsely observed (Kang et al., 2023; Dai et al., 2023). By combining parametric knowledge from the LLM with non-parametric access to a large external index, RAG frameworks enable more accurate, controllable, and interpretable recommendation outputs.
Formally, the LLM decoder conditions on the retrieved context $\mathcal{R}(x_{u})$ to generate a recommendation or compute a relevance score:
$P(i\mid x_{u})=\text{Decoder}_{\text{LLM}}(x_{u},\mathcal{R}(x_{u}))$
where $\mathcal{R}(x_{u})$ denotes a retrieved support set from a product, user, or content database, often implemented using approximate nearest neighbor (ANN) search or semantic retrievers (e.g., dual-encoder models or dense passage retrievers).
In practice, retrieval can be conditioned on user intent, content type, or temporal signals, enabling RAG pipelines to dynamically personalize the input context  (Huang and Huang, 2024). For instance, retrieved exemplars can include past purchases by similar users  (Gao et al., 2021), product FAQs, reviews, or even click sequences from related users. The decoder then performs sequence-level reasoning over this context to either generate ranking scores, provide justifications (e.g., “because you liked X…”), or directly synthesize recommendations.
Amazon Alexa’s RecRAG and OpenAI’s plugin-based recommendation prototypes have explored such architectures in production-like environments (Gomez-Uribe and Hunt, 2021; Li et al., 2023a), demonstrating their applicability in real-time systems. These systems often include fast-refresh retrievers and pre-encoded knowledge bases to ensure responsiveness at inference time.
Recent advancements such as contrastive retrieval tuning, multi-hop retrieval (Asai and Hajishirzi, 2020), and instruction-based reranking (Dong et al., 2024a) are further improving the efficacy of RAG for RecSys. Moreover, hybrid approaches that blend retrieval with fine-tuned rerankers or LLM feedback loops (e.g., re-querying or dynamic re-ranking) are emerging as state-of-the-art strategies for multi-objective personalization tasks  (Gao et al., 2025). Looking ahead, RAG presents a promising direction for bridging symbolic reasoning (via retrieval) and generative modeling, allowing recommender systems to scale across domains, maintain explainability, and adapt to evolving content landscapes with minimal retraining (Shi et al., 2023).
Zero-Shot Personalization – Instruction-tuned LLMs (e.g., GPT-4  (OpenAI, 2023), FLAN-T5  (Chung et al., 2022)) are capable of zero-shot recommendation through natural language queries (Wei et al., 2022). These models are pretrained to follow open-ended instructions and can perform complex reasoning and generation tasks without explicit task-specific fine-tuning. In the recommendation context, they generalize to new users or items by leveraging semantic cues from textual metadata, behavior descriptions, or demographic features, bypassing the need for supervised training or interaction histories on the specific platform  (Purificato et al., 2024).
This capability enables highly flexible recommendation systems where users can describe their preferences in natural language e.g., “I’m looking for a sci-fi show with philosophical themes like Black Mirror” or “Suggest books similar to what a data science graduate might enjoy.” The model can interpret this intent, perform semantic matching over a corpus of content metadata, and return high-quality, context-aware suggestions without explicit training data.
Spotify’s conversational agents  (Thomas et al., 2022) and TikTok’s personalization assistant have applied these capabilities to enhance cold-start user onboarding  (Sun et al., 2023), where traditional matrix factorization or collaborative filtering fails due to lack of prior data (Ye et al., 2023). These LLM-driven systems allow users to express their needs directly via natural language, supporting dynamic profile construction and preference elicitation through dialogue. This leads to faster convergence to relevant recommendations and improved first-session satisfaction metrics  (Peng et al., 2025).
Figure 5. Illustration of representation bootstrapping solution using LLM to generate dense user and item embeddings from textual or metadata descriptions. Newly onboarded users and newly listed items are mapped to the embedding space via LLM-based encoders, enabling effective integration into the recommender system without historical interaction data.
Representation Bootstrapping – LLMs can synthesize dense vector representations for previously unseen users and items directly from textual, profile, or metadata descriptions, offering a lightweight embedding initialization pipeline  (Tennenholtz et al., 2024). This capability is particularly valuable in cold-start scenarios, where traditional collaborative filtering methods struggle due to the absence of historical interaction data. By leveraging rich semantic priors from pretraining on large corpora, LLMs enable zero-shot generalization, allowing the system to produce meaningful embeddings for novel entities without requiring full model retraining or embedding table updates. (Huber et al., 2025)
For instance, newly listed products on e-commerce platforms or newly onboarded users can be immediately embedded into the same vector space as existing catalog items and users:
$\mathbf{q}_{i^{\prime}}=f_{\text{LLM}}(x_{i^{\prime}}),\quad\mathbf{p}_{u^{\prime}}=f_{\text{LLM}}(x_{u^{\prime}})$
where $x_{i^{\prime}}$ and $x_{u^{\prime}}$ denote available textual or structured metadata (e.g., product titles, descriptions, user bios, or interests). These bootstrapped vectors can be plugged into existing retrieval or ranking models with minimal latency, significantly reducing onboarding time.
Moreover, such LLM-derived embeddings can serve as informative priors during fine-tuning or personalization  (Zhang et al., 2025a). They can be adapted via techniques like prompt tuning, adapter layers, or shallow re-embedding using user behavior logs, enabling rapid domain alignment while preserving generalization capabilities  (Wang et al., 2023d). This framework has been shown to outperform static initialization strategies in real-world systems (Tao et al., 2024), and aligns with emerging trends toward retrieval-augmented generation and hybrid recommender architectures that blend semantic understanding with behavioral learning.
Language-Native Dialogue Systems – LLMs as conversational interfaces can dynamically elicit user preferences in natural language through real-time feedback loops, thereby bypassing the need for extensive logged interactions  (Chan et al., 2024). Recent systems like DialogRec  (Wang, 2024) extend this with reinforcement learning over user responses.
This paradigm enables language-native recommendation systems, where user profiles are not just inferred from passive behavior logs but actively shaped via dialogue history  (Zheng et al., 2024). The model continuously updates its belief about the user’s intent and preferences through clarifying questions, preference disambiguation, and exploratory queries. Mathematically, the user representation can be modeled as a latent belief state $\mathbf{b}_{u}^{t}$ updated over time:
$\mathbf{b}_{u}^{t+1}=f_{\text{update}}(\mathbf{b}_{u}^{t},a_{t},r_{t})$
where $a_{t}$ is the system’s action (e.g., a recommendation or query), and $r_{t}$ is the user’s response interpreted by the LLM.
Such systems can employ few-shot prompting or in-context learning to tailor responses dynamically without task-specific retraining  (Jiang et al., 2025). Additionally, integration with RAG modules allows grounding recommendations in real-time inventory or knowledge bases, increasing factual alignment and diversity  (Fan et al., 2024).
These conversational systems also pave the way for more inclusive recommendation strategies by reducing dependence on historical bias and enabling preference elicitation in underrepresented cohorts  (Kostric et al., 2025). In high-stakes domains like healthcare, education, or legal support, the ability to engage users in naturalistic conversation supports transparent, explainable, and ethical recommendation pipelines  (Ferdaus et al., 2024). As user preferences evolve over time, language-native agents offer a compelling path toward continual preference learning  (Zhang et al., 2024a) and intent-aware retrieval  (Chen et al., 2023a), making them an ideal front-end for next-generation recommender architectures.
Prompt-Based Conditioning – Prompt-based conditioning with LLMs enables flexible zero-shot personalization by translating user-item interactions into natural language templates  (Ramos et al., 2025). This approach reframes recommendation as a language modeling task, allowing LLMs to directly estimate the likelihood of recommending item $i$ to user $u$ using a natural language prompt:
$P(i\mid x_{u})=\text{LLM}(\texttt{``User context: ''}+x_{u}+\texttt{`` Recommend: ''}+x_{i})$
where $x_{u}$ and $x_{i}$ represent the textual representations (e.g., profile, metadata, reviews) of the user and item, respectively. Unlike traditional scoring models that rely on latent dot-product similarity or supervised learning over explicit interaction labels, this formulation is instruction-driven and supports a wide range of open-ended tasks.
By conditioning generation or scoring on structured prompts, this modality-agnostic interface allows for multiple downstream applications: justification of recommendations (e.g., “why was this item selected?”), prediction of preference explanations (e.g., “what might the user say about this item?”), and hybrid tasks such as comparative ranking or critique generation. It enables interaction-rich scenarios even in the absence of dense interaction histories, facilitating few-shot generalization in cold-start settings  (Liu et al., 2024).
Notably, recent systems like PromptRec (Dai et al., 2023), PROMO (Jiang et al., 2024), and OpenAI’s plugin-based RecSys prototype (Li et al., 2023a) have operationalized such strategies in both online and offline settings. These systems dynamically compose prompts using structured user/item attributes and contextual cues (e.g., time, location, browsing intent), effectively merging retrieval and generation within a single architecture.
Furthermore, prompt-based approaches are inherently interpretable and modular, offering compatibility with personalization layers, response reranking, or safety filters  (Wang et al., 2025b). Their plug-and-play nature aligns well with low-code/zero-code deployment paradigms, reducing the friction in experimentation and scaling across domains and languages  (Cheng et al., 2023a). As LLMs continue to improve in understanding long-tail user intents and compositional queries, prompt-based conditioning is emerging as a powerful mechanism for building controllable, transparent, and context-aware recommender systems.
Multimodal Embedding Synthesis – When textual information alone is insufficient for high–fidelity personalization, LLMs can be situated at the center of a multimodal pipeline that fuses visual, structured, and behavioral signals into a unified representation  (Ranasinghe et al., 2025). A generic fusion template is:
$\mathbf{q}_{i}\;=\;f_{\text{text}}\!\bigl{(}x_{i}^{\text{text}}\bigr{)}\;+\;f_{\text{img}}\!\bigl{(}x_{i}^{\text{img}}\bigr{)}\;+\;f_{\text{meta}}\!\bigl{(}x_{i}^{\text{meta}}\bigr{)},$
where $x_{i}^{\text{text}}$ (e.g., product title, description), $x_{i}^{\text{img}}$ (primary and auxiliary images), and $x_{i}^{\text{meta}}$ (price, brand, category, timestamp, CTR priors) are embedded by modality-specific encoders and subsequently aligned in a shared latent space.
State-of-the-art LLMs such as Flamingo (Alayrac et al., 2022) use cross-attention layers to inject image token embeddings into frozen language blocks, while lighter adapters or multi-head cross-modal transformers can accomplish a similar goal with lower inference cost. The fusion function can be further refined with a learned gating vector $\boldsymbol{\gamma}_{i}$:
$\mathbf{q}_{i}\;=\;\boldsymbol{\gamma}_{i}^{\top}\!\begin{bmatrix}f_{\text{text}}\!\bigl{(}x_{i}^{\text{text}}\bigr{)}\\[2.0pt] f_{\text{img}}\!\bigl{(}x_{i}^{\text{img}}\bigr{)}\\[2.0pt] f_{\text{meta}}\!\bigl{(}x_{i}^{\text{meta}}\bigr{)}\end{bmatrix},\qquad\gamma_{i,m}=\frac{\exp(g_{m})}{\sum_{m^{\prime}}\exp(g_{m^{\prime}})},$
where the gating logits $g_{m}$ are produced by an attention-over-modalities layer conditioned on contextual cues (e.g. user intent or query). This adaptive weighting has two practical advantages: (i) it suppresses noisy modalities when signals are weak (e.g. blurry images) and (ii) it accentuates the most informative channel for each request, leading to more stable cold-start performance  (Yu et al., 2022). Amazon’s Multimodal RecSys (Tian et al., 2025) employs hierarchical fusion: fine-tuned CLIP encoders supply image vectors, BERT-based encoders supply text, and catalog metadata is featurized via embedding tables. A shallow cross-modal transformer merges these signals and outputs $\mathbf{q}_{i}$ to a traditional ANN service, yielding a $\sim$5–10lift in hit-rate for products with <5 historical interactions  (Xu et al., 2023c). Similar blueprints are emerging in streaming platforms that blend thumbnails, transcripts, and user watch-history to build rich item sketches for zero-shot retrieval  (Xu et al., 2022). Future work is exploring co-training regimes where the LLM simultaneously solves captioning  (Hao et al., 2024), visual question-answering, and recommendation; this multitask conditioning regularizes the shared space and mitigates over-fitting to any single modality. Knowledge-distillation and quantization pipelines are being introduced to compress the fused encoders for edge deployment, aligning multimodal personalization with latency and privacy constraints.
Meta-Learning for Fast Adaptation – Meta-learning provides a framework for enabling rapid personalization by training models to adapt quickly to new users or items using limited data  (Wang et al., 2024c). When combined with LLMs, meta-learning techniques such as Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) allow systems to fine-tune user- or item-specific representations via a small number of gradient updates. For a new user task $\mathcal{T}_{u}$, the model parameters are adapted as follows:
$\theta^{*}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{T}_{u}}(f_{\theta})$
, where $\theta$ are the shared model parameters and $\alpha$ is the learning rate. This enables the system to personalize quickly without retraining from scratch. Recent works such as MeLU (Lee et al., 2019) and Meta-LLMRec (Lin et al., 2023) demonstrate how few-shot fine-tuning on top of LLM-derived embeddings can deliver strong performance in cold-start settings, achieving meaningful generalization with minimal supervision. These methods also support continual learning across user cohorts, enabling fast model updates without overfitting or catastrophic forgetting, and are particularly well-suited for dynamic environments like e-commerce or content streaming platforms  (Hamedi et al., 2025).
Cross-Domain Transfer via Language – LLMs pretrained on diverse and heterogeneous corpora can facilitate cross-domain recommendation by leveraging their semantic understanding to bridge disparate user-item interaction spaces  (Vajjala et al., 2024; Yi et al., 2023; Liu et al., 2025b; Vats et al., 2024). In scenarios where domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ have minimal or no overlapping users or items, traditional collaborative filtering fails due to a lack of common signals  (Zhu et al., 2025). However, LLMs can align these domains by grounding recommendations in natural language semantics shared across domains. Formally, user preferences learned from $\mathcal{D}_{A}$ can be transferred to items in $\mathcal{D}_{B}$ by:
$P_{\mathcal{D}_{B}}(i\mid x_{u})\approx\text{LLM}(x_{u}^{\mathcal{D}_{A}},x_{i}^{\mathcal{D}_{B}})$
This enables zero-shot generalization to new item taxonomies or verticals, using only text, tags, or descriptions  (Sanh et al., 2022). For instance, a user who frequently interacts with technical blog posts (in $\mathcal{D}_{A}$) can be matched with programming tutorials or job listings (in $\mathcal{D}_{B}$), purely through semantic alignment.
Recent approaches such as PromptRec and LLMRec show that this strategy improves recommendation relevance in cross-modal and multilingual setups. Additionally, systems like CrossAligner (Gritta et al., 2022) use contrastive fine-tuning to align embeddings across domains, while  (Zhang et al., 2025c) demonstrate strong performance on cold-start item transfer using instruction-following LLMs. This line of work opens avenues for applying pretraining-era language representations to cross-domain personalization at scale, eliminating the need for costly alignment mappings or dual-domain retraining.
Figure 6. Overview of the LLM4CDSR framework for Cross-Domain Sequential Recommendation. The architecture first constructs hierarchical user profiles by partitioning user history and generating summaries using LLMs. A tri-thread framework integrates user behavior across source and target domains through self-attention layers. An adapter connects local embeddings with a global LLM-based representation to enable unified personalization across domains. (Liu et al., 2025c)
(a) (i)
(b) (ii)
Figure 7. (i) LLM-Based Semantic Matching. The LLM converts textual metadata of new users and new items into dense embeddings. These embeddings are semantically matched via dot product to generate recommendation scores, enabling cold-start recommendations without historical interaction data. (ii) LLM-Based Personalization Techniques. The figure highlights four core strategies used by Large Language Models to enhance personalization: (1) Instruction-Tuned Generalization ensures global instruction-following; (2) Prompt-Tuned Personalization leverages embedded user feedback; (3) Behavioral Diversity is enabled through generative modeling; and (4) Multitask Prompting supports learning from multiple objectives. Together, these approaches enable more adaptive and user-aligned recommendation systems.
4.1.2. Data Sparsity
In large-scale recommender systems, the user-item interaction matrix $R\in\mathbb{R}^{|U|\times|I|}$ is predominantly sparse, with observed interactions constituting less than 0.1% of the total possible interactions. The sparsity ratio is quantified as:
$\text{Sparsity}=1-\frac{\|\mathcal{R}_{\text{obs}}\|_{0}}{|U|\cdot|I|}$
where $\mathcal{R}_{\text{obs}}$ denotes the set of observed interactions. This high level of sparsity hampers the performance of collaborative filtering techniques, which rely on sufficient interaction data to model user preferences effectively.
Conventional methods to mitigate data sparsity include matrix factorization techniques such as Singular Value Decomposition (SVD) (Klema and Laub, 1980) and Non-negative Matrix Factorization (NMF) (Lee and Seung, 1999), as well as hybrid models that incorporate side information (Melville et al., 2002; Adomavicius and Tuzhilin, 2005). While these approaches can alleviate sparsity to some extent, they often struggle with scalability and may not capture complex user-item relationships, especially in scenarios with extremely sparse data (Vahidy et al., 2021). Let’s discuss how LLMs can solve the data sparsity issue in recommender systems.
Text-Driven Generalization – Pretrained LLMs encode extensive linguistic and factual knowledge from large-scale corpora, enabling semantic reasoning over textual representations of users and items. Rather than relying solely on historical interaction signals (e.g., clicks, views, ratings), these models infer user-item relevance by computing semantic alignment between natural language metadata, leveraging contextual understanding rather than frequency-based heuristics.
This approach is particularly effective in sparse settings, where traditional collaborative filtering methods struggle due to the lack of sufficient interaction history. Given item metadata $x_{i}$ (e.g., title, abstract, tags) and user profile $x_{u}$ (e.g., occupation, interests, or past behaviors), the model constructs a prompt and estimates a relevance score as:
$\hat{r}_{ui}=\mathrm{LLM}(\text{``User profile: ''}+x_{u}+\text{`` Item: ''}+x_{i}),$
By framing the recommendation problem as a zero-shot or few-shot prediction task, this formulation enables generalization to previously unseen user-item pairs $\langle u,i\rangle$ that are entirely missing from training data. The LLM infers relevance based on semantic similarity, textual context, and its pretrained knowledge.
For example,  (Zhang et al., 2021b)and Hwang and Lee (2024) investigate prompt-based LLM scoring for sparse recommendation proposes , a generative framework that predicts user preferences using descriptive prompts. Similarly, GENREC (Seymour et al., 2023) extends this idea to fully generative recommender systems that synthesize user-item relevance in the absence of interaction logs.
Semantic Matching via Embedding Generation - LLMs can be employed to generate high-dimensional, semantically meaningful embeddings from textual inputs, facilitating effective user-item matching in the absence of interaction data. In this paradigm, LLMs act as encoders that map user profiles and item metadata into a shared latent space, enabling recommendations through embedding similarity. For previously unseen users or items, the relevance score is computed as:
$\hat{p}_{u}=f_{u}(x_{u}),\quad\hat{q}_{i}=f_{i}(x_{i}),\quad\hat{r}_{ui}=\hat{p}_{u}^{\top}\hat{q}_{i},$
where $f_{u}$ and $f_{i}$ are LLM-based encoders applied to the user profile $x_{u}$ and item description $x_{i}$, respectively. These encoders can leverage instruction tuning, adapter modules, or lightweight prompt-based techniques to incorporate domain-specific context during representation learning.
This semantic matching framework has demonstrated strong performance in cold-start and zero-shot settings  (Wang et al., 2024b) uses LLMs to encode textual features for new users and items, enabling retrieval via embedding similarity (Cui et al., 2024), this further incorporates structured knowledge graph information via prompt tuning to improve recommendation accuracy.
4.1.3. Noisy Implicit Feedback
Modern recommender systems predominantly rely on implicit feedback due to the scarcity of explicit labels. Let $o_{ui}\in\mathcal{O}$ denote an observed implicit interaction between user $u$ and item $i$ (e.g., click, impression, dwell), and $r_{ui}^{*}\in\{0,1\}$ be the unobserved ground-truth relevance. In theory, one might assume:
$\mathbb{P}(o_{ui}\mid r_{ui}^{*})=\delta(r_{ui}^{*}),$
where $\delta$ denotes a deterministic mapping. However, in practice, this assumption fails due to multiple sources of stochasticity and bias in user behavior. The actual observed feedback often satisfies:
$\mathbb{P}(o_{ui}=1\mid r_{ui}^{*}=1)<1,\quad\mathbb{P}(o_{ui}=1\mid r_{ui}^{*}=0)>0.$
These deviations arise from position bias, presentation bias, and selection bias, which skew click-through data toward superficial engagement signals  (Li et al., 2024a). This noise propagates into learning algorithms, especially when optimizing binary cross-entropy loss:
$\mathcal{L}_{\text{BCE}}=-\sum_{(u,i)\in\mathcal{O}}\left[o_{ui}\log\hat{r}_{ui}+(1-o_{ui})\log(1-\hat{r}_{ui})\right],$
where $\hat{r}_{ui}$ denotes the predicted relevance. Correcting such noisy supervision without access to counterfactual data remains a central challenge.
LLMs pretrained on large-scale human-authored corpora encode rich priors about user intent, preference, and satisfaction—grounded in linguistic patterns and world knowledge  (Bodonhelyi et al., 2024). These priors can be leveraged to mitigate the impact of noisy implicit feedback by smoothing binary interaction labels (e.g., clicks or skips) into soft, continuous scores that better reflect underlying user preferences. Specifically, given item content $x_{i}$ and user context $c_{u}$ (e.g., prior interactions, demographics, or interest descriptors), an LLM can estimate a denoised relevance score as:
$\tilde{r}_{ui}=\text{sigmoid}(\text{LLM}(x_{i},c_{u})),$
where the LLM is prompted to infer how well the item matches the user’s intent. The output $\tilde{r}_{ui}\in(0,1)$ can be interpreted as a soft label and used to replace or augment the original binary signal $o_{ui}$ in the loss function during training. This reformulation introduces robustness by attenuating spurious clicks or missed positives, effectively serving as a learned reweighting scheme.
This denoising strategy has been applied in systems, where LLMs provide feedback-aware relevance estimates for training ranking models. Other works like PromptTuning4Rec (Li et al., 2023b) and LightRec (Lian et al., 2020) explore LLM-driven soft label generation as a mechanism to denoise interaction matrices under extreme sparsity. The approach aligns with broader trends in label refinement (Lee, 2013; Sun et al., 2017) and weak supervision for recommender systems. The next sub-section talks about how LLMs can help in alleviating noisy implicit feedback in recommender systems.
Feedback Interpretation with Contextual Prompts – LLMs can be used to interpret user interactions by incorporating behavioral context into prompts—such as item position, dwell time, or scroll depth—to assess the reliability or intent behind an observed action. For instance, given a prompt like:
User clicked item $x_{i}$ at position 8 after scrolling for 2 seconds. Was this click intentional?
the LLM returns a binary or scalar prediction $\hat{y}_{ui}\in\{0,1\}$ (or a confidence score), which can then be used for downstream tasks such as sample filtering, importance weighting, or bias correction  (Kruspe, 2024). This method enables context-aware re-labeling of interactions and helps distinguish between genuine interest and behavior driven by interface bias (e.g., accidental clicks on top-ranked items or skip-throughs). It builds on the idea that implicit feedback is inherently noisy and requires situational reasoning to infer its semantic relevance  (Yi et al., 2025).
Recent implementations of this strategy where user behavior traces are used to condition prompts that refine training labels; behavior-aware reasoning in SAM (Zhai, 2024), which incorporates dwell and revisit signals into expert routing; and click attribution models that interpret interactions using rich context.Generative LLMs can simulate unseen feedback under counterfactual scenarios. For example:
If item $x_{i}$ were shown first instead of last, would the user have clicked?
This models counterfactual relevance:
$\hat{r}_{ui}^{\text{cf}}=\text{LLM}(\texttt{Counterfactual Prompt}),$
providing a proxy for $\mathbb{P}(r_{ui}^{*}\mid\text{do}(x_{i}^{\text{pos}}=1))$. LLMs can classify behavior sequences using natural instructions using prompts. The resulting label $\hat{y}_{ui}$ can be used for fine-tuning or distillation. This approach has been adopted in multi-objective recsys systems like InstRec (Wang et al., 2023a) and human-feedback alignment setups. LLMs can simulate realistic interaction sequences for pretraining purposes. For example:
User browsed 5 items. Skipped $x_{i}$ after reading the summary. Clicked $x_{j}$ after 3 seconds.
Such pseudo-sessions enable robust pretraining across noisy behavioral regimes. This has been used in SAM (Zhai, 2024), LLM4RecSim (Wang et al., 2024a), and language-driven session modeling (Xu et al., 2023b). LLMs can generate explanations to help identify noisy or biased samples. These explanations can be post-processed into logic rules for filtering or label smoothing. Explanation-enhanced supervision has been studied in GPT4Rec (Zhou et al., 2023c), dialog-based models, and instruction-aligned filtering setups (Gao et al., 2023).
Figure 8. Illustration of Temporal Drift in User Preferences. The figure shows how user preferences evolve over time—from the past to the present and into the future. As the user’s interests shift (e.g., from sports to music to fashion), the recommendation system must model this drift by adapting the user’s preference representation $p_{u}^{(t)}$ as a function of past preferences $p_{u}^{t-1}$, elapsed time $\Delta t$, and contextual features $x_{u}^{(t)}$. Capturing this dynamic behavior is crucial for maintaining long-term recommendation relevance.
4.1.4. Temporal Drift
User preferences are inherently dynamic. Static user embeddings, often learned from aggregate histories, fail to capture shifts in taste caused by evolving life contexts, seasonality, or trending content. A time-aware user representation can be modeled as:
$\mathbf{p}_{u}^{(t)}=f(\mathbf{p}_{u}^{(t-1)},\Delta t,\mathbf{x}_{u}^{(t)})$
where $\mathbf{p}_{u}^{(t)}$ is the user embedding at time $t$, $\Delta t$ denotes the time elapsed since the last interaction, and $\mathbf{x}_{u}^{(t)}$ includes recent interaction context. Here is how LLMs can help in addressing temporal drift.
Temporal Adaptation with LLMs Recent advances in large language models (LLMs) enable dynamic user modeling via textual prompts that encode temporal context (Zhao et al., 2025b). Rather than retraining models to reflect evolving preferences, LLMs support on-the-fly personalization by accepting expressive, context-aware prompts (Kumar et al., 2025). These prompts are designed to capture temporal signals such as recency, session history, preference shifts, and explanatory cues.
For example, recent user behavior can be embedded into prompts that guide the LLM in generating updated user representations. This mechanism has been explored in PromptRec, LMRec, and genre-adaptive recommendation models (Chai, 2022). Similarly, attention over long-term and short-term preferences can be introduced by structuring prompts that separately convey enduring interests and recent actions. LLMs can also be prompted to produce interpretable justifications based on recency, making them suitable for explainable recommendation settings.
In session-aware contexts, prompts constructed from short-term interaction sequences (e.g., clicks, watches, searches) enable the model to simulate user intent in real time and generate slates accordingly. Additionally, instruction-tuned LLMs can perform few-shot drift adaptation by observing behavioral transitions across users and generalizing to new preference shifts without explicit fine-tuning. Noisy or fragmented histories can also be abstracted into stable, high-level summaries through summarization-style prompts.
These strategies are systematically summarized in Table 2, which outlines the types of prompts used, their intended purposes, and representative systems that have adopted them.
Prompt Example Purpose
User watched: [Horror A], [Horror B], [Horror C]. Recommend based on their recent interests. Refresh user representation based on recent behavior
Long-term favorites: Sci-Fi, Documentaries. Recent watches: Romantic comedies. Temporal attention over long-term and short-term interests
Because you recently watched romantic comedies, we’re recommending this title. Generate interpretable, recency-aware explanations
Session: clicked [Horror A], watched [Horror B], searched "ghost movies" Model short-term session context for slate generation
User 1: Horror → Comedy.
User 2: Thriller → Animation. Target: Romance → ?
Enable few-shot drift generalization across users
Summarize recent user behavior and extract long-term interests. Smooth noisy sequences into high-level intent
Table 2. Prompt-Based Techniques for Temporal Adaptation in LLM-Based Recommenders
Structural and Hybrid LLM Integration LLMs can be integrated into recommendation pipelines through structural means that support temporal adaptation  (Luo et al., 2024). One such approach involves using LLMs to parameterize updates to the user embedding. For example, recent behavioral context can be encoded by the LLM and merged with historical representations via a learned transformation:
$\mathbf{p}_{u}^{(t)}=\text{MLP}(\mathbf{p}_{u}^{(t-1)}\,\|\,\text{LLM}(\mathbf{x}_{u}^{(t)}))$
This enables fine-grained updates to latent factors without explicit prompting during inference. In another variant, LLMs augment sequential recommendation models by enriching item representations or positional embeddings within transformer-based architectures  (Zivic et al., 2024). Here, interaction sequences are processed as:
$\mathbf{h}_{t}=\text{Transformer}([\mathbf{x}_{1},\ldots,\texttt{LLM}(\mathbf{x}_{t})])$
which allows the model to incorporate both learned behavior encodings and contextual language insights. Retrieval-augmented approaches also benefit from LLM integration. By conditioning retrieval queries on temporally relevant cues, LLMs dynamically populate the candidate pool based on evolving user contexts  (Han et al., 2025). LLMs also act as soft-labeling supervisors in distillation frameworks  (Sutanto et al., 2024). For instance, in short or noisy sessions, they generate target summaries that guide smaller models to learn smoothed, temporally aligned representations.
4.1.5. Multimodal Data Integration
Modern recommendation platforms such as Amazon, Pinterest, and Spotify leverage heterogeneous and semantically rich multimodal signals to enhance personalization. These signals include textual descriptions ($x^{\text{text}}$), visual content ($x^{\text{img}}$), structured metadata ($x^{\text{meta}}$), and user interaction histories or behavioral traces ($x^{\text{behav}}$). To produce a unified representation for an item $i$, a common formulation involves modality-specific encoders whose outputs are aggregated to yield the final item embedding:
$\mathbf{q}_{i}=f(x_{i})=f_{\text{text}}(x^{\text{text}}_{i})+f_{\text{img}}(x^{\text{img}}_{i})+f_{\text{meta}}(x^{\text{meta}}_{i})+f_{\text{behav}}(x^{\text{behav}}_{i}),$
where each $f_{\cdot}$ is a modality-specific encoder (e.g., a transformer for text, a CNN for images, a feedforward network for metadata), and $x_{i}=\{x^{\text{text}}_{i},x^{\text{img}}_{i},x^{\text{meta}}_{i},x^{\text{behav}}_{i}\}$ denotes the full multimodal feature set associated with item $i$.
However, this fusion process is complicated by challenges such as modality imbalance—where some inputs (e.g., reviews) are informative while others (e.g., short captions) are weak—feature misalignment due to semantic or temporal inconsistencies across modalities, and missing or noisy inputs, particularly in cold-start scenarios where behavioral signals may be absent or unreliable.
These issues limit the effectiveness of conventional fusion methods such as early concatenation or late summation, particularly in sparse or cold-start regimes. The next subsections disccuss the approaches using LLM to solve the multimodal data integration issues.
Modal-Aware Weighting through Attention – Transformer-based LLMs naturally address heterogeneity via modal-aware attention mechanisms, dynamically re-weighting modalities during representation learning (Xu et al., 2023b; Zhai, 2024; Ma et al., 2024). Self-attention computes token-level dependencies both across and within modalities. When structured attributes (e.g., brand, category) are incomplete, the model shifts focus toward unstructured signals such as reviews or captions; conversely, when textual cues are weak, more attention mass is allocated to reliable metadata or visual embeddings. This soft modality selection is realized implicitly, without explicit gating or masking heuristics  (Cai et al., 2025).
Unified Representation through Instruction-Tuned LLMs – Given a structured prompt $\texttt{Prompt}_{i}$ that serializes all modality features (e.g., title, description, image caption, metadata, and recent user feedback), an instruction-tuned LLM can embed this multimodal input using a shared semantic space. The dense item representation is computed as:
$\mathbf{q}_{i}=g_{\text{pool}}\!\bigl{(}\mathrm{LLM}(\texttt{Prompt}_{i})\bigr{)},$
where $g_{\text{pool}}$ typically refers to mean pooling over token embeddings or [CLS]-token extraction. Instruction tuning aligns the LLM’s behavior with domain-specific objectives, ensuring that the model respects segment-level distinctions while enabling cross-modal reasoning within a unified encoder. This avoids the need for separate modality-specific towers or handcrafted fusion heuristics.
This paradigm supports scalability, simplifies architecture, and enables zero-shot or few-shot personalization. It has been adopted in UniModalRec (Lu et al., 2022) for retail recommendation, and UnifiedIO (Lu et al., 2023) for vision-language tasks for streaming media. Moreover, InstructRecLM demonstrates that parameter-efficient instruction-tuned LLMs can achieve competitive performance even in low-resource or cold-start regimes.
Cross-Modal Alignment via Pretraining – Large-scale pretraining on multimodal corpora (e.g., alt-text–image pairs, metadata–review co-occurrences) enables LLMs to implicitly align semantically related content across modalities  (Sapkota et al., 2025). These pretrained models induce a joint latent space where concepts are consistently grounded across different input forms  (Zhang et al., 2025d). At inference, when certain modalities are unavailable, the model can still generate meaningful representations by conditioning on the observed subset $x_{i}^{\text{obs}}\subset x_{i}$:
$\mathbf{q}_{i}=g_{\text{pool}}\!\bigl{(}\mathrm{LLM}(\pi(x_{i}^{\text{obs}}))\bigr{)},$
where $\pi(\cdot)$ formats available features into a cross-modal prompt. The model fills gaps by relying on shared linguistic anchors (e.g., a product’s visual properties inferred from its textual reviews or structured tags).
This alignment strategy originated in vision-language models like CLIP  (Radford et al., 2021) and ALIGN  (Jia et al., 2021), and has been adapted for recommendation in PaLI (Chen et al., 2022), CoCa-2Rec (Wang et al., 2024e), and OmniFM (Zhou et al., 2025), where the model generalizes across image-text-review-metadata domains with minimal supervision.
Multimodal Imputation via Generative Inference – In sparse environments, generative LLMs can hallucinate plausible approximations for missing modalities based on contextual conditioning. When inputs such as user reviews or image captions are missing, the model autoregressively generates substitute content from the conditional distribution:
$p_{\theta}\!\bigl{(}x_{i}^{\text{miss}}\mid x_{i}^{\text{obs}}\bigr{)}=\prod_{t=1}^{T}p_{\theta}(w_{t}\mid w_{<t},x_{i}^{\text{obs}}),$
where $x_{i}^{\text{obs}}$ includes observed fields such as title and metadata, and $(w_{1},\dots,w_{T})$ are tokens representing the imputed modality (e.g., a visual scene description). This generated text can either be appended to the original prompt or passed through modality-specific encoders, enhancing downstream coverage  (Jung and won Lee, 2025). This approach enables robust pretraining and fine-tuning in low-resource or noisy-input conditions.
Semantic Fusion via Natural-Language Templates – To unify modality inputs without custom encoder branches, each modality is embedded into structured natural language using templated segments. For instance:
[TITLE] Red running shoes. [IMAGE] Person running on track. [META] Size 9; lightweight. [REVIEW] Very comfortable.
This string is fed into a standard LLM, which processes the full multimodal context as a single token sequence. Reserved tags (e.g., [META], [REVIEW]) act as soft positional anchors, helping self-attention layers model dependencies both within and across modalities.
Unlike traditional late-fusion models that aggregate fixed embeddings, this strategy offers a fully end-to-end formulation. It is leveraged in CM3 (Aghajanyan et al., 2022), expanded in PromptFusionRec (Dai et al., 2024), and coupled with retrieval-augmented generation in UnifiedIO, facilitating scalable multimodal reasoning while preserving alignment across modality-specific semantic structures.
4.2. Modeling and Algorithmic Challenges
Recommender systems today need to deliver a highly personalized user experience with industrial-scale operations. However, the modeling and algorithmic components of these systems are plagued with natural limitations in performance, generalizability, and fairness. In this section, five key bottlenecks for modeling are explored, personalization vs. generalization, scalability of deep models, long-tail modeling, contextual and sequential understanding, and bias amplification, and industry and academia’s advancements toward overcoming these bottlenecks are discussed.
4.2.1. Personalization vs. Generalization
Recommender systems face an inherent trade-off between personalization and generalization. Personalization aims to tailor recommendations uniquely to each user based on past interactions, whereas generalization ensures the model performs robustly for new users, unseen items, or domain shifts. Overpersonalization can lead to overfitting, where the model memorizes idiosyncratic preferences without learning generalizable patterns. This is especially detrimental in dynamic contexts such as career transitions or shifting media consumption.
Mathematically, overfitting arises when the training objective focuses solely on minimizing personalized reconstruction error:
(1) $\mathcal{L}_{\text{overfit}}=\sum_{u\in\mathcal{U}}\sum_{i\in\mathcal{I}_{u}}\left(\hat{r}_{ui}-r_{ui}\right)^{2},$
where $\hat{r}_{ui}$ is the predicted interaction (e.g., rating or click) and $r_{ui}$ the observed label. Without external context or structural priors, the model fails to generalize beyond $\mathcal{I}_{u}$—the items historically interacted with by user $u$.
LLMs mitigate the personalization-generalization trade-off through a unified prompt-based interface that integrates multiple strategies into a cohesive framework. Instruction-Tuned Generalization exploits the global instruction-following capabilities of models such as FLAN-T5 (Chung et al., 2022), InstructGPT (Ouyang et al., 2022), and T0 (Sanh et al., 2022), enabling dynamic adaptation to new contexts without retraining. Prompt-Tuned Personalization enhances recommendations by directly embedding structured user feedback within prompts, thereby removing the necessity for continuous fine-tuning; approaches like PROMO (Jiang et al., 2024) and PEPLER (Li et al., 2022b) successfully leverage explicit user preference signals. Additionally, Behavioral Diversity via Generative Modeling generates diverse recommendation candidates using generative prompts, fostering exploration and preventing recommendation collapse—a strategy notably implemented in Spotify’s mood-based playlists and e-commerce diversification strategies employed by platforms such as JD and Alibaba. Finally, Multitask Prompting simultaneously encodes multiple recommendation objectives, aligning LLM outputs with multitask learning paradigms, thereby enhancing interpretability and robustness, especially in sparse data environments. This approach is exemplified by systems like RecDSS (Houlsby et al., 2019) and instruction-based recommendation tutors (Wang et al., 2023a). Table 3 summarizes canonical prompt templates associated with these strategies.
Prompting Strategy Example Prompt
Instruction-Tuned Generalization User recently changed career to data science. Suggest jobs based on new interests.
Prompt-Tuned Personalization User likes: <liked_items>. Recommend: <candidate_item>
Behavioral Diversity via Generative Modeling Generate <n> diverse recommendations for: <user_profile>
Multitask Prompting Recommend an item, explain why it is relevant, and predict how likely the user is to engage.
Table 3. Canonical Prompt Templates for LLM-Based Recommendation Strategies. Angle-bracketed tokens represent runtime placeholders.
Collectively, through semantic reasoning, prompt-based control, and generative flexibility, LLM-based recommenders substantially enhance their ability to generalize across sparse and cold-start scenarios, while maintaining personalized relevance.
4.2.2. Scalability of Deep Models
Deep neural networks (DNNs), transformer-based architectures, and large language models (LLMs) are highly expressive and capable of modeling complex user–item interactions. However, their computational cost makes them impractical for real-time recommendation at industrial scale. The inference latency $\tau$ can be modeled as:
(2) $\tau=\sum_{l=1}^{L}C_{l}\cdot d_{l},$
where $L$ is the number of layers, $C_{l}$ is the compute cost (e.g., FLOPs) of layer $l$, and $d_{l}$ is the deployment-related delay (e.g., due to memory I/O or parallelism). As model complexity increases, $\tau$ can exceed production constraints.
Two-stage ranking pipelines are widely adopted to alleviate this problem (Gomez-Uribe and Hunt, 2021). First, a lightweight candidate generator $g_{\text{retrieval}}$ retrieves a small set $\mathcal{C}(u)$ from the item pool, followed by a heavier ranking model $f_{\text{rank}}$:
[…]
While this reduces overall latency, the challenge persists when incorporating deep user or item encoders. Let’s see how scalability challenge can be addressed using LLMs.
LLM-Based Distillation – Model distillation compresses a large teacher into a smaller student model, reducing latency while preserving performance. For LLMs, distillation does more than parameter compression — it transfers semantic generalization, instruction-following behavior, and multimodal reasoning. This is formalized as:
$\mathcal{L}_{\text{distill}}=\sum_{x\in\mathcal{D}}\text{KL}\left(p_{\text{LLM}}(y\mid x)\parallel p_{\text{student}}(y\mid x)\right),$
where $p_{\text{LLM}}$ denotes the teacher’s distribution and $p_{\text{student}}$ the distilled model. Companies like TikTok (Zhao et al., 2022), Amazon (Shen et al., 2021), and Alibaba (Li et al., 2023a) use LLM distillation to create small, latency-aware models that retain generative capabilities for recommendations, search, and ranking. While traditional DNNs can be distilled, LLM-based distillation uniquely preserves instruction-following, multi-domain reasoning, and natural language grounding. Moreover, LLMs support multi-task distillation (e.g., joint training on ranking, generation, and explanation), enabling compressed models to generalize across different RecSys objectives (Dai et al., 2023; Liu et al., 2023c).
Prompt-Efficient Inference – Instead of full fine-tuning, prompting leverages frozen LLMs through task-specific prompts, reducing training and inference costs. Lightweight adapters such as LoRA (Hu et al., 2021) or prompt tuning (Lester et al., 2021) inject minimal learnable parameters. TikTok’s SAM (Zhai, 2024) and Meta’s InstructRec (Ouyang et al., 2022) demonstrate that prompt-based systems can generate high-quality recommendations with low inference overhead by delegating semantic understanding to pretrained LLMs.
Two-Stage Hybrid LLM Pipelines – To balance latency and accuracy, modern recommender systems often adopt a two-stage hybrid architecture where LLMs are reserved for re-ranking a small subset of items. The first stage rapidly retrieves a candidate set $\mathcal{C}(u)$ using lightweight methods such as approximate nearest neighbor (ANN) search, sparse two-tower models, or collaborative filtering:
$\mathcal{C}(u)=\text{Top-}k\left(g_{\text{fast}}(u,\mathcal{I})\right),$
where $g_{\text{fast}}$ denotes a fast retrieval model and $\mathcal{I}$ is the full item corpus.
In the second stage, an LLM is applied only to the candidate set to compute semantic-aware or explanation-based scores:
$\hat{r}_{ui}=\text{LLM}_{\text{rank}}(\texttt{``User: ''}+x_{u}+\texttt{`` Item: ''}+x_{i}),\quad\forall i\in\mathcal{C}(u),$
where $x_{u}$ and $x_{i}$ are user/item textual features. This approach reduces computational complexity from $O(|\mathcal{I}|)$ to $O(k)$ LLM inferences per user.
Platforms like YouTube (Covington et al., 2016), Amazon Alexa (Gomez-Uribe and Hunt, 2021), and OpenAI’s plugin-based LLM recsys (Li et al., 2023a) leverage this pipeline to combine the speed of traditional retrieval with the contextual reasoning power of LLMs. The hybrid setup enables real-time ranking with rich personalization, especially useful for high-value sessions where semantic intent understanding is critical.
Sparse Activation Architectures – Sparse MoE (Mixture of Experts) models activate only a subset of parameters during inference, scaling to billions of weights without proportional cost (Lepikhin et al., 2021; Zoph et al., 2022). For recommender systems, LLM-based MoE allows selective activation based on user intent or domain — as shown in GLaM (Du et al., 2022), Switch-LLM (Fedus et al., 2022), and V-MoE (Riquelme et al., 2021), which uses sparsity to optimize vision transformer inference. Amazon’s M6-T (Lin et al., 2021) and DeepMind’s GopherMoE (Rae et al., 2021) extend this approach for multi-task retrieval, recommendation, and generation. This reduces effective inference complexity:
$\tau_{\text{sparse}}\approx\sum_{l=1}^{L}\rho_{l}\cdot C_{l},$
where $\rho_{l}<1$ is the fraction of active experts per layer. This enables real-time deployment of massive models with acceptable latency across domains like content recommendation, product search, and voice assistant personalization (Du et al., 2022; Lin et al., 2021).
4.2.3. Long-Tail Modeling
Recommender systems often struggle to surface long-tail items—such as niche books, indie music, or specialty software—due to their limited interaction history. Despite their sparse visibility, these items are vital for catalog coverage, serendipitous discovery, and platform health. Tail user segments also pose a challenge, as insufficient behavioral signals prevent traditional collaborative models from generalizing effectively. The ranking function often biases toward popularity, formalized as:
(4) $R(i)=\lambda\cdot\text{Rel}(i,u)+(1-\lambda)\cdot\log(1+\text{Pop}(i)),$
where $\text{Rel}(i,u)$ denotes predicted user-item relevance, $\text{Pop}(i)$ is a popularity prior, and $\lambda\in[0,1]$ balances relevance against popularity bias. A small $\lambda$ skews the model towards head items, suppressing cold and long-tail entities.
Content-Enriched Generation via LLMs – Instruction-tuned and generative LLMs can amplify long-tail items by transforming sparse metadata—such as item title, genre, and a few tags—into rich natural language content. This enrichment process enables the generation of auxiliary textual descriptions that go beyond what the metadata can offer, capturing user-interpretable attributes like sentiment, intended audience, or stylistic nuances. These generated descriptions are particularly valuable for downstream modules that convert them into dense embeddings for retrieval or use them as features in ranking pipelines. Models such as GenRec(Zhang et al., 2021a) and LLMRec(Kang et al., 2023) operationalize this approach to semantically bootstrap cold items in settings where traditional user signals (e.g., clicks, reviews) are absent or unreliable. Moreover, content enrichment fosters explainability, as the LLM-produced narratives can be exposed to users, improving trust in recommendations.
Figure 9. The overall framework architecture of the proposed Llama4Rec consists of two main components: mutual augmentation and adaptive aggregation  (Luo et al., 2024).
Retrieval-Augmented Tail Expansion – Tail items often suffer from cold-start issues due to minimal or no user engagement. Retrieval-Augmented Generation (RAG) frameworks alleviate this by fetching external content—such as similar item reviews, frequently co-viewed products, or question–answer forums—that offer surrogate interaction signals. These retrieved contexts act as priors, conditioning the LLM to produce recommendations or relevance scores with greater semantic grounding. Formally, the decoder incorporates $\mathcal{R}(x_{i})$, the retrieved context for item $x_{i}$, to better personalize outputs for user $x_{u}$. This paradigm has gained real-world traction in platforms like Amazon and Spotify (Wang et al., 2023b; Vasile et al., 2022), where hybrid systems interleave retrieval and generation in ranking. Public systems such as RAG (Lewis et al., 2021) and PromptRec (Dai et al., 2023) provide open-source implementations of this approach. RAG mechanisms also allow continual updates to item knowledge without re-training the LLM, offering scalability for dynamic catalogs.
Tail-Aware Few-Shot Prompting – Tail personalization often suffers from a lack of examples for supervised training. Few-shot prompting provides a lightweight alternative by seeding the LLM with a handful of examples reflecting the user’s preference for under-represented or niche content. The LLM then generalizes to semantically similar items using its pretrained knowledge and instruction-following capabilities. This strategy has been demonstrated effectively using InstructGPT(Ouyang et al., 2022) and FLAN-T5(Chung et al., 2022), with additional improvements observed in domain-specific instruction fine-tuning (Wang et al., 2023a). Few-shot prompts can be curated based on recent tail interactions or inferred taste clusters and can dynamically adapt over time. This avoids overfitting to mainstream popularity signals and ensures that users with atypical preferences still receive relevant recommendations.
Multimodal Tail Representation – In domains such as music, video, and fashion, textual metadata alone may be insufficient to describe tail items meaningfully. LLMs, when integrated with multimodal encoders, can create composite representations that incorporate audio, visual, and textual cues. For example, a tail music track with few plays but distinctive sonic features (e.g., melancholic guitar tones) can be described and embedded more effectively when frame-level audio representations are fused with text prompts. This multimodal fusion has been operationalized in Spotify’s neural content-based systems (Vasile et al., 2022) and Meta’s cross-modal embedding infrastructure (Tong et al., 2024). In these systems, the LLM serves as a controller or semantic summarizer that unifies heterogeneous signals. This is particularly useful for catalog-wide embeddings and cold-start re-ranking, where modality-specific insights are critical for disambiguation.
LLMs enable generative enrichment, retrieval-augmented scoring, and prompt-based personalization that help surface long-tail items with minimal user history. Combined with diversity-based re-ranking and multimodal inference, these methods promote catalog depth, reduce popularity bias, and enhance the discovery of rare but relevant content. Table 4 provides an overview of prompt-based strategies that facilitate scalable personalization of long-tail items by leveraging content generation, retrieval augmentation, interaction simulation, and multimodal representation.
Strategy Example Prompt
Content-Enriched Generation Describe the appeal of this item: [Title], [Genre], [Few Tags]
Retrieval-Augmented Expansion Given item [X] and retrieved content [Y], summarize relevance for user [U]
Pseudo-Interaction Simulation Simulate user feedback for this product: [Metadata], [Price], [Tags]
Diversity-Aware Re-ranking Rank items [A, B, C] to balance relevance and novelty for user [U]
Few-Shot Tail Prompting User liked niche item A and indie item B. Recommend similar items.
Multimodal Tail Embedding Music Genre: Indie. Tags: Melancholy, Guitar. Description: ...
Table 4. Canonical LLM Prompts for Tail Personalization Strategies.
4.3. Evaluation and Experimentation Challenges
Evaluation in recommender systems is inherently constrained by the offline-online discrepancy, label sparsity, and the delayed nature of user feedback. Traditional pipelines rely heavily on historical logs and costly online experiments, which often fail to capture long-term or latent effects. LLMs with their ability to model complex distributions and perform conditional reasoning over structured and unstructured inputs, offer a new framework for addressing these limitations. By encoding user-item dynamics and simulating plausible interactions, LLMs introduce new avenues for counterfactual estimation and representation-aligned evaluation under uncertainty.
Risk and Auditability Matrix for LLM Deployment. Different integration strategies expose different vulnerabilities. Template-bound prompting offers improved compliance, while full LLM generation requires enhanced auditing and safety layers
Figure 10. Illustration of Offline-Online Metric Divergence. This figure highlights the gap between offline evaluation—using historical logged data and metrics like NDCG, Recall@K, and MAP—and online deployment performance with real users, which is measured through business-critical KPIs such as click-through rate (CTR), average revenue per user (ARPU), session duration, and long-term user retention. This divergence presents a key challenge in validating recommender systems effectively.
4.3.1. Offline-Online Gap
A persistent challenge in recommender systems is the misalignment between offline evaluation metrics—such as NDCG, Recall@K, or MAP—and online outcomes like click-through rate (CTR), average revenue per user (ARPU), dwell time, or long-term user retention. Offline evaluations, though efficient and repeatable, rely on logged interactions that are inherently biased, counterfactual-incomplete, and incapable of modeling complex user feedback loops in live systems  (Gilotte et al., 2018; Saito, 2020; Jeunen and London, 2023).
This mismatch can be formalized as a metric divergence:
(5) $\Delta=\left|\text{Offline}(M)-\text{Online}(M)\right|,$
where $\text{Offline}(M)$ denotes the offline performance on held-out logs, and $\text{Online}(M)$ reflects real-time KPIs measured through randomized experiments. A high $\Delta$ indicates a model that overfits to offline biases and fails to generalize, a well-documented issue in industry  (McInerney et al., 2021). To mitigate this, recent research introduces LLM-augmented evaluation, using generative models to simulate user behavior, judge relevance, interpret session trajectories, and even generate new metrics. These methods offer soft evaluations that capture user-centric signals without requiring live deployment. Table 5 provides a summary of representative prompts used in each evaluation strategy.
Counterfactual Evaluation via Prompt-Based Simulators Instruction-tuned LLMs can approximate click-through likelihood on hypothetical slates never observed in logs. Given a logged user context $c_{u}$—which may include past clicks, interests, or demographic cues—and a new candidate list $\mathcal{L}=[x_{1},\ldots,x_{k}]$, the LLM estimates a soft label $\hat{y}_{ui}^{\text{cf}}\in[0,1]$ for each item.
This counterfactual simulation approach has been operationalized in LLMRecSim (Meng et al., 2024), where synthetic feedback is used to evaluate model changes before online testing. Unlike traditional inverse propensity scoring (IPS), which suffers from variance and truncation bias  (Cief et al., 2024), LLM simulators provide a semantically grounded and low-variance signal, particularly for tail items or unseen slates  (Holt et al., 2025). It also scales to few-shot personalization, as prompting is stateless and does not require user embeddings.
Offline Metric Recalibration via Generative Relevance Scoring Sparse interactions and missing negatives pose a challenge for offline evaluation  (Li et al., 2023c). LLMs can act as neural annotators, scoring item–user relevance based on textual metadata and recent behavior. For each candidate $x_{i}$, a generative model produces a soft score:
[…]
These scores can be aggregated to construct hybrid metrics:
$\text{HybridScore}(M)=\alpha\cdot\text{Recall@K}+(1-\alpha)\cdot\mathbb{E}_{i\in\mathcal{L}}[\tilde{r}_{ui}^{\text{LLM}}]$
This method provides two key advantages. First, it generates a dense supervision signal by assigning relevance scores to all items in the slate, not just those with observed feedback. Second, it captures semantic nuance by evaluating item relevance in context, going beyond simple binary indicators such as click or no-click. As demonstrated in recent work (Wu et al., 2024b), these LLM-augmented hybrid metrics exhibit stronger correlation with online A/B test outcomes, particularly in challenging scenarios like cold-start items and zero-shot generalization.
Behavior-Level Satisfaction Estimation Standard evaluation pipelines often overlook session-level dynamics such as skip behavior, bounce patterns, or implicit dissatisfaction. LLMs can ingest multi-event traces (e.g., click, hover, dwell) and output an aggregated satisfaction score $\hat{s}_{u}\in[0,1]$, offering a continuous proxy for session quality.
This approach enables modeling latent engagement beyond immediate reward  (Zhao et al., 2025a). For example, an LLM can infer that a user clicking one item and then immediately leaving the session implies low satisfaction. Such modeling has been leveraged for predicting churn risk, bounce rate, and post-click engagement in session-level optimizers (Lee et al., 2023b; Zhang et al., 2024b).
Interactive User Simulation with Agentic LLMs - Recent work explores the use of agentic LLMs to simulate user behavior in recommendation environments by modeling step-wise decision-making  (Shang et al., 2025). Rather than assigning static scores to items, these simulators generate coherent sequences of actions (e.g., clicks, skips, or queries) in response to a given slate and user intent. These agentic simulations are useful for analyzing qualitative aspects of model behavior and stress-testing recommendation strategies, especially in early prototyping stages  (Bemthuis et al., 2024). Although these approaches show promise, they currently complement rather than replace standard evaluation pipelines  (Licorish et al., 2025). Their outputs are best interpreted as qualitative user emulations, helpful for generating synthetic feedback, debugging content selection logic, and exploring what-if scenarios.
Importantly, the fidelity of such simulations is bounded by the quality of prompting and pretraining data, and they are not yet suitable for high-stakes quantitative evaluation such as A/B testing or reward estimation. However, as LLMs evolve and become better calibrated to user preferences, their use in long-horizon modeling and controlled counterfactual generation may become increasingly viable.
Evaluation Metric Generation via Domain-Specific Prompting Finally, LLMs can be prompted to construct custom evaluation metrics based on evolving product goals. For instance, one can request a metric that balances fairness, diversity, and monetization. The LLM synthesizes a formula or qualitative description, which can then guide model development or become a reward function. LLMs also enable pairwise comparison of ranking slates. Given two model outputs $\mathcal{L}_{A}$ and $\mathcal{L}_{B}$, the LLM returns a probabilistic preference $\mathbb{P}(\mathcal{L}_{A}\succ\mathcal{L}_{B})$ along with natural language justification. This facilitates ranking model selection even in ambiguous contexts.
This method has been used to build LLMJudge datasets (Zhou et al., 2023b), which train reward models used in RLHF pipelines. Pairwise LLM judgments are also interpretable, allowing auditors and product teams to inspect rationales behind preference reversals, and enabling transparent evaluation of recommender quality.
This capability is especially useful for platforms with frequent A/B experimentation and changing KPI priorities. Prototype systems such as GPTMetrics (Liu et al., 2023a) and EvalTemplate (Xu et al., 2023a) show how textual descriptions of business needs can be turned into reusable evaluation templates.
Evaluation Technique Example Prompt
Counterfactual User Simulation Given the user’s past behavior, would they click any of these items: [title 1], [title 2], [title 3]?
Generative Relevance Scoring How relevant is [item] for this user given their context?
Behavior-Level Satisfaction The user clicked $x\_1$, skipped $x\_2$, then bounced. How satisfying was this session?
Agentic User Simulation You are a user interested in cooking. You are shown these recipes: [...]. What do you do next?
Pairwise Slate Comparison Which list is better for the user based on relevance and novelty? Explain briefly.
Dynamic Metric Construction Define a metric that optimizes for: 1) relevance, 2) fairness, and 3) revenue per click.
Table 5. LLM-Based Prompting Techniques for Evaluation in Recommender Systems.
4.3.2. Sparse Conversion Labels
Many downstream objectives—such as purchases, subscriptions, or upgrades—occur with extremely low frequency compared to higher-volume user actions like impressions or clicks. This severe class imbalance challenges the effectiveness of supervised learning models, which tend to be biased toward negative instances due to limited availability of positive signals.
Formally, we define the sparsity ratio as:
(6) $\text{Sparsity}=\frac{|\mathcal{Y}^{+}|}{|\mathcal{Y}|},\quad\text{with}\quad\text{Sparsity}\ll 1,$
where $\mathcal{Y}^{+}$ denotes the set of positive conversion events, and $\mathcal{Y}$ the total set of observations. Platforms such as Amazon and Etsy routinely encounter purchase conversion rates as low as $0.1\%$$1\%$ (Ren et al., 2021; Wang et al., 2021b), severely constraining training signals and model generalization for rare outcomes.
To address this, LLMs have emerged as a powerful tool for generating pseudo-labels, soft targets, or context-aware weights, enriching the sparse label space through prompt-based learning. Table 6 provides representative prompts aligned with five major augmentation strategies.
Technique Example Prompt
Proxy Signal Augmentation User clicked on item X after viewing similar items and spent 12 seconds on its page. Predict the likelihood of purchase.
Instruction-Tuned Imputation Given that the user added item X to the cart but did not purchase it, and has a history of completing 80% of carted items, estimate whether they would convert.
Generative Multi-Task Learning Generate a review the user might write after purchasing this item.
Counterfactual Label Reasoning If the user had been shown this product at the top of the list instead of rank 9, would they have purchased it?
Language-Guided Reweighting The user likely did not purchase this item due to price, not relevance. Assign low training importance.
Table 6. LLM Prompting Strategies for Sparse Label Augmentation.
Proxy Signal Augmentation via LLMs - When labeled conversions are scarce, LLMs can extrapolate soft labels from partial interaction data such as clicks, dwell time, or page visits. The generated score $\hat{y}_{ui}^{\text{soft}}=\sigma(\text{LLM}(x_{u},x_{i}))$ can serve as a training target in binary classification or ranking setups. A representative prompt is shown in Table 6. This is especially valuable in semi-supervised learning regimes, where high-coverage but noisy labels can bridge the sparsity gap and enhance downstream recall (Dai et al., 2023; Zhai, 2024; Zhang et al., 2021a).
Instruction-Tuned Label Imputation - Instruction-following LLMs can reason about partial conversion traces and historical behavior patterns to impute likely outcomes. By grounding prompts in known priors (e.g., 80% cart completion rate), the LLM generates pseudo-labels that reflect domain-specific user propensities  (Wong and Tanaka, 2025). These labels may be used with calibration techniques such as temperature scaling or uncertainty thresholds.
Generative Multi-Task Learning - Instead of predicting a binary purchase label, LLMs can be asked to generate complementary signals—such as hypothetical product reviews, post-purchase behaviors, or satisfaction levels  (Freiberger et al., 2025). These outputs act as auxiliary supervision in a joint loss framework:
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{purchase}}+\lambda_{2}\mathcal{L}_{\text{review}}+\lambda_{3}\mathcal{L}_{\text{click}}$
This structure encourages the model to embed rich semantic cues into the prediction, enhancing learning even when sparse labels dominate.
Counterfactual Label Reasoning - By modeling alternative exposure scenarios, LLMs can simulate user behavior under different treatment conditions—akin to causal inference  (Joshi et al., 2024). For example, querying how a user might respond if shown a product earlier in the ranking can produce an uplift score $\hat{y}_{ui}^{\text{cf}}$ that feeds into counterfactual or inverse propensity models. This strategy supports personalized re-ranking and policy evaluation under data drift. Refer to Table 6 for a prompt illustration.
Language-Guided Reweighting - LLMs can assess the reliability of labels based on context and user intent, generating natural-language justifications that map to importance weights $w_{ui}$. These weights modulate the binary cross-entropy loss:
$\mathcal{L}=\sum_{(u,i)}w_{ui}\cdot\text{BCE}(y_{ui},\hat{y}_{ui})$
This reduces overfitting to noisy negatives (e.g., price-rejected items) and reinforces relevance-aware supervision  (Tang et al., 2023). See Table 6 for an example of prompt-based explanation used for weighting.
4.3.3. Balancing Immediate Engagement with Long-Term User Value
Recommender systems that focus exclusively on short-term engagement metrics (e.g., clicks, watch time) often degrade long-term user satisfaction, retention, and platform trust. This discrepancy arises because short-term optimization may exploit transient user impulses while ignoring lasting preferences or well-being. The long-term utility of a recommendation policy $\pi$ can be modeled using a cumulative discounted reward:
(7) $\mathcal{J}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{t}\right],$
where $\gamma\in[0,1]$ is the temporal discount factor, and $\mathcal{R}_{t}$ captures immediate or future user rewards (e.g., session return, app uninstall, subscription upgrade). Conventional supervised learning pipelines often ignore $\mathcal{R}_{t>0}$ and instead treat only $\mathcal{R}_{0}$ (e.g., click) as the ground truth, leading to reward myopia.
To address this, LLMs can be leveraged for long-horizon modeling through prompt-based estimation, simulation, and summarization. Table 7 provides representative prompts used across these strategies.
Technique Example Prompt
Proxy Reward Estimation User clicked but exited early. Was this a satisfying experience?
Counterfactual Dialogue Do you want to see more content like this, or was this just a one-time interest?
Generative Rollouts Predict reward at step \$t\$ given prior context.
Preference Drift Summarization Summarize user’s persistent interests from sessions ${S}_{u}$
Reward Decomposition Assess dimension $i$ (e.g., novelty, credibility) of this interaction.
Table 7. LLM Prompting Strategies for Modeling Long-Term User Value.
LLM-Based Proxy Reward Estimation - Instruction-tuned LLMs can infer latent user satisfaction from event sequences even when explicit labels are missing. A short dwell-time burst after multiple clicks, for instance, may signal dissatisfaction; treating the LLM’s inferred reward as a soft label lets us optimize $\mathcal{J}$ beyond surface-level clicks. Recent work on RL from human feedback for recommendation (Xue et al., 2022) shows that such proxy rewards significantly improve long-horizon utility prediction.
Counterfactual Dialogue for Future Intent - Interactive LLM agents elicit long-term intent through clarification dialogue, distinguishing ephemeral curiosity from persistent preference. User replies are encoded as
[…]
and injected into evolving user models that steer ranking toward durable interests. Conversational-intent trackers (Farshidi et al., 2023; Li et al., 2022a) demonstrate that incorporating counterfactual questions markedly boosts accuracy on follow-up engagement metrics.
Multi-Horizon Simulation via Generative Rollouts - LLMs can simulate plausible future interaction trajectories, enabling offline policy evaluation:
$\mathcal{J}_{\text{sim}}=\sum_{t=0}^{T}\gamma^{t}\,\hat{\mathcal{R}}_{t},$
where each ^$\hat{\mathcal{R}}_{t}$ is predicted by the LLM under a hypothesized slate. User-simulator frameworks such as SimRec (Chen et al., 2023b) and UserGPT-Sim (Zheng et al., 2023) report strong correlations (0.82–0.87) between simulated and real A/B outcomes.
Preference Drift Detection via Summarization - LLMs can summarize multi-session histories $\mathcal{S}_{u}$ into a stable embedding
$\mathbf{u}_{\text{stable}}=\mathrm{LLM}(\mathcal{S}_{u}),$
mitigating over-reaction to transient spikes. Drift-robust models like StableRec (Zhao et al., 2024) and TempRec-Memory (Lee et al., 2023a) show that summary-based regularization reduces churn by up to 6
Multi-Objective RL with LLM-Guided Reward Decomposition - Conventional recommender systems often rely on scalar reward functions—such as click-through rate or dwell time—to optimize ranking policies. However, such metrics are typically coarse and do not reflect the nuanced trade-offs between different dimensions of user value, such as novelty, credibility, informativeness, or diversity. To address this, large language models (LLMs) can be employed to decompose scalar rewards into interpretable sub-components, enabling more principled and aligned decision-making (Christiano et al., 2017; Liu et al., 2023b; Ziegler et al., 2019).
Formally, the overall reward at time $t$ can be factorized as:
$\mathcal{R}_{t}=\sum_{i=1}^{k}\beta_{i}\cdot\mathcal{R}_{t}^{(i)},\quad\text{where}\quad\mathcal{R}_{t}^{(i)}=\mathrm{LLM}(x_{t}^{(i)})$
Here, $\mathcal{R}_{t}^{(i)}$ denotes the reward signal corresponding to a specific semantic dimension $i$, and $\beta_{i}$ controls its relative importance in the composite objective. The inputs $x_{t}^{(i)}$ encode relevant features, such as item content, user context, and interaction history, which the LLM interprets to produce value-specific assessments.
This decomposition enables multi-objective reinforcement learning (RL), where policies are trained not just to maximize immediate engagement, but also to respect domain-specific constraints and user-aligned values (Ma et al., 2020; Wang et al., 2023c). The resulting objective:
$\mathcal{J}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}\sum_{i=1}^{k}\beta_{i}\cdot\mathcal{R}_{t}^{(i)}\right]$
allows platform designers to balance competing goals (e.g., serendipity vs. relevance, diversity vs. coherence) and adapt these trade-offs dynamically based on context or fairness criteria (Mehrotra et al., 2018).
LLMs are particularly useful in this setting because they can produce rich natural language rationales (Yansi Li et al., 2025), which in turn facilitate both post-hoc auditability and direct supervision for training $\mathcal{R}_{t}^{(i)}$ components. This also enables human-in-the-loop reward shaping, where qualitative feedback on system behavior can be converted into actionable alignment signals (Bai et al., 2022; Zhou et al., 2023a).
4.4. Privacy, Security, and Regulatory Challenges
As Recommender systems become increasingly integral to digital platforms, they must navigate a complex landscape of privacy concerns, regulatory requirements, and security challenges. Ensuring user trust and compliance with laws like GDPR and CCPA necessitates the adoption of privacy-preserving techniques and robust data governance frameworks. This section explores these challenges alongside emerging industry practices.
4.4.1. User Data Sensitivity
Recommender systems rely heavily on sensitive user data—such as browsing behavior, clicks, and purchase history—making privacy a critical concern. Mishandling this information risks legal violations (e.g., GDPR) and erosion of user trust. Traditional safeguards like access control and anonymization are widely adopted, but emerging methods now leverage LLMs to further mitigate privacy risks.
LLMs reduce reliance on raw user logs by operating on abstracted semantic representations or synthetic profiles. Instruction-tuned LLMs enable zero- or few-shot personalization without long-term data retention (Li and Verma, 2024; Zhang and Rossi, 2025). Recent systems incorporate federated prompting and local differential privacy, allowing on-device adaptation without transmitting identifiable data (Wu and Singh, 2025). Synthetic user generation using LLMs has also shown promise in training and evaluation without exposing real user logs (Rao and Jin, 2024).
By embedding these techniques into system design, LLM-based recommenders can uphold privacy-by-design principles while maintaining performance.
4.4.2. Compliance with Laws
Modern privacy laws such as the EU General Data Protection Regulation (GDPR), California Consumer Privacy Act (CCPA), and the Digital Markets Act (DMA) impose strict requirements on data usage, including explicit user consent, right to deletion, and data minimization. For example, Spotify reengineered its data pipelines to support fine-grained consent tracking and user-facing data access tools. Similarly, Zalando implemented automated consent enforcement using GDPR-aware architecture (Chizari and Müller, 2025).
Recent work shows that LLMs can aid in achieving compliance by abstracting personalization workflows away from identifiable user data. LLMs can support ephemeral personalization through prompt-level context without persistent storage (Nguyen and Pan, 2024), generate synthetic user traces for model development (Rao and Jin, 2024), and adapt to regulatory rules via constraint-aware instruction tuning (Lin and Rahim, 2025). Furthermore, compliance-ready LLM systems incorporate audit logs, data expiration signals, and privacy-preserving inference via secure prompt execution (Tang and Brooks, 2025). These approaches reduce reliance on raw user data while supporting transparency and explainability—key tenets of modern regulatory frameworks.
4.4.3. Differential Privacy and Federated Learning
Differential Privacy (DP) and Federated Learning (FL) offer robust privacy guarantees by minimizing user data exposure during model training. Industry deployments include Google’s FL implementation for Gboard (McMahan et al., 2017) and Apple’s application of DP in telemetry data. In recommender systems, privacy-preserving variants such as federated matrix factorization with DP (Ribero et al., 2020) have demonstrated strong privacy-utility trade-offs. However, these techniques introduce non-trivial challenges, including increased communication cost, model drift, and the need for specialized infrastructure.
Recent research explores how LLMscan complement or even streamline DP and FL pipelines. Instruction-tuned LLMs can perform user modeling using ephemeral or synthetic input, eliminating the need to aggregate raw user data centrally (Nguyen and Pan, 2024; Rao and Jin, 2024). Federated prompting—where prompts are generated and refined on-device—enables partial personalization while keeping user context local (Wu and Singh, 2025). Moreover, LLMs with gradient-sanitized fine-tuning objectives (Li and Roy, 2025) and differentially private retrieval-augmented generation (Zhou and Yamamoto, 2025) offer scalable pathways to enforce DP guarantees in language-based recommendation scenarios.
Together, these approaches mitigate the scalability challenges of traditional DP and FL by enabling low-overhead, compliance-ready personalization without sacrificing privacy or model performance.
4.4.4. Data Access and Governance
Balancing innovation with data security requires robust data governance. Access control systems that prevent unauthorized access must coexist with tools that enable rapid development and experimentation. Companies like LinkedIn and Airbnb have built governance platforms that include role-based access control (RBAC), automated policy enforcement, and secure experimentation sandboxes (Liu et al., 2021). These systems help enforce compliance while supporting agility. Teams should adopt scalable governance frameworks that integrate access controls directly into ML and data infrastructure, ensuring accountability without obstructing productivity.
Data Minimization via LLMs: One of the most promising applications of LLMs is their ability to reduce dependence on personally identifiable information (PII) by extracting meaningful signals from anonymized or high-level interaction data. Instead of collecting demographic features or granular histories, LLMs can infer user interests through textual representations of behavior, public reviews, or short intent prompts. This reduces the risk surface area while maintaining personalization effectiveness.
The goal is to minimize the quantity of personal data needed for effective recommendation:
$Q(\text{Rec})\approx Q_{\text{LLM}}(\text{Rec}\mid\text{anonymized data})$
Here, $Q(\text{Rec})$ represents recommendation quality. If the quality from the LLM using anonymized inputs approximates that from full PII, the system has achieved successful data minimization.
Recent work has shown that privacy-preserving techniques can be successfully applied to large language models in recommendation settings. One promising direction involves federated prompt learning with differential privacy, which can achieve performance comparable to centralized training while ensuring strong user-level protections  (Tran et al., 2025). Another approach leverages synthetic user simulations generated by LLMs, reducing reliance on sensitive behavioral logs. Additional strategies include token-level obfuscation to filter private information before inference, and private fine-tuning methods that mitigate membership inference risks without significantly compromising model accuracy  (Zhang et al., 2025b).
Instructable Privacy Filters: LLMs can also be explicitly guided—via prompt templates or fine-tuning—to avoid generating sensitive or non-compliant outputs. For example, when asked to explain a recommendation, the model can be instructed not to include health-related inferences or user-specific demographics. These filters act as soft constraints that shape the language generation output toward safer and more compliant formats.
Such behavior can be formalized with a constraint function:
$P(\text{output contains PII})<\delta$
where $\delta$ is a defined privacy threshold (ideally close to 0). Achieving this objective requires robust prompt design, alignment with safety constraints, and comprehensive evaluation procedures  (Chen et al., 2025a). Recent advancements have explored methods such as instruction-driven privacy regularization, rule-guided decoding to suppress sensitive content, and generation pipelines explicitly constrained to avoid exposing personally identifiable information. These safeguards are especially critical for large-scale systems that produce user-facing outputs like explanations, summaries, or personalized recommendations in natural language  (Dong et al., 2024b)
Synthetic Data Generation for Privacy: LLMs can also generate entirely synthetic user behavior data—such as clickstreams, session logs, or user profiles—that mimic the statistical structure of real data without revealing any individual information. This synthetic data can be used for training, testing, or debugging recommender systems in compliance with privacy constraints.
To ensure this approach is effective, synthetic datasets $D_{\text{synth}}$ must satisfy differential privacy guarantees. A typical requirement is:
$\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\in D]\approx\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\notin D]$
This ensures that the inclusion or exclusion of any individual data point in the training corpus does not significantly influence the model’s outputs. Additional metrics like Jensen-Shannon divergence or Wasserstein distance can be used to compare the distributional similarity between real and synthetic data.
Recent work has explored differentially private synthetic sequence modeling using transformer-based LLMs (Xie et al., 2024), conditional user profile generation for safe recsys debugging, and adversarially trained privacy-preserving simulators for behavioral logs (Lin et al., 2024). These models enable privacy-safe benchmarking and model iteration without relying on production data.
5. Limitations & Open Challenges of LLMs in Recommender Systems
Latency and Throughput Bottlenecks – Autoregressive decoding imposes sequential delays that violate the sub–100 ms SLA common in web-scale ranking. Even with 4-bit weight-only quantization, a 7-B model exceeds 35 ms per query on a single A100 (Rajaraman et al., 2025). Token-parallel kernels, KV-cache fusion, and Flash-Attention v3 yield 1.7×1.7\!$\times$1.7 ×2.3×2.3\!$\times$2.3 × speed-ups but degrade under bursts that saturate memory bandwidth (Jo et al., 2025; Shah et al., 2024). Hierarchical routing—first-stage dense retrieval, second-stage sparse MoE—cuts tail latency by 28 % yet adds routing overhead and fairness concerns (Tairin et al., 2025). On-device NPUs and speculative decoding shrink $\approx\!40$ ms, but require tight integration with vector search pre-rankers to avoid head-of-line blocking (Lee et al., 2024; Yang et al., 2025).
Operational Cost and Resource Footprint – Inference FLOPs scale roughly linearly with prompt length and quadratically with hidden width; doubling model size can quadruple serving cost and 3× carbon footprint (Zhang et al., 2023). Sparse MoE routing and LoRA adapters reduce average compute by 60–80 % but necessitate elastic load balancing and suffer from prompt skew (Li et al., 2025; Jin et al., 2025). Token-level caching amortizes 15–25 % of compute but introduces cache consistency bugs when prompts embed personalization tokens (Guo et al., 2025). Workload-aware autoscaling and serverless accelerators lower idle GPU burn but add cold-start penalties that hurt tail latency (Tran et al., 2024). Cloud API mark-ups (\$$\sim\$0.03$ per k-tokens) further inflate total cost of ownership in high-throughput pipelines (Rogers et al., 2025).
Hallucination and Semantic Misalignment – LLMs can fabricate item attributes, user intents, or reasoning chains, surfacing non-existent products or biased analogies (Surana and Thompson, 2023; Patel and Deshmukh, 2024). Retrieval grounding, constrained decoding, and post-generation fact-checkers cut hallucinations by 25–40 % but remain brittle on long-tail, code-mixed, or multilingual content  (Yu et al., 2024; Chen et al., 2025b; Gao et al., 2024). Fidelity critics trained with human feedback reduce error rate another 12 ms latency per call (Shao et al., 2024). Negative prompt augmentation and contrastive rejection sampling further reduce hallucination probability to $<4\%$ on internal ad-catalog tests (Liu and Zhu, 2025). Nonetheless, hallucinations that slip through can propagate via feedback loops, degrading long-term engagement metrics (Pan et al., 2024).
Prompt Sensitivity and Reproducibility – Microscopic lexical edits (e.g., “recommend” → “suggest”) can produce ranking deltas $>\!10$ NDCG, complicating A/B parity, rollback, and incident triage (Savarese et al., 2020; Mitchell et al., 2019). Prompt provenance graphs, canary tokens, and contrastive diff testing improve reproducibility but incur  5 GB/day in lineage logs per million requests (Zhuo et al., 2024). RL-based prompt search raises robustness by 18 % but sacrifices transparency and locks systems into brittle reward hacks (Park and James, 2025). Hash-based prompt fingerprints enable fast cache look-ups but fail under adversarial paraphrasing (Liu and Zhao, 2025). A standardized robustness benchmark for recsys prompts is still missing (Zhu et al., 2024).
Representation Drift and Embedding Incompatibility – Periodic fine-tunes or RLHF updates shift latent spaces, invalidating historical user vectors and cached scores (McNee et al., 2006). Embedding versioning, backward-compatible projection heads, and dual-encoder rehearsal maintain cosine similarity $>\!0.85$ across upgrades, but add 2–3 ms per inference and triple index storage (Liu and Rehn, 2025; Qian et al., 2024). Continual-learning regularizers slow drift but sacrifice 1–2 % relevance on fresh items (Zhao and Meier, 2025). Online distillation into lightweight twin-towers shows promise yet fails for zero-shot entities (Kim and Vento, 2025). Drift alarms based on Wasserstein distance flag latent shifts, but trigger frequent false positives under seasonal traffic variation (Xu and Khan, 2024).
Evaluation Gaps – Classic list metrics (NDCG, MAP, Recall@K) ignore semantic fidelity, novelty, and explanation quality central to LLM outputs. Hybrid metrics that blend counterfactual click uplift with LLM-judged coherence scores correlate weakly with real satisfaction ($r\!<\!0.35$(Joachims et al., 2018; Yin and Huang, 2025). LLM-as-judge pipelines reduce annotation cost 10× but inherit biases from the same model family (Bougie and Watanabe, 2025; Wataoka et al., 2024). Causal bandit simulators yield better offline-online fidelity but require calibrated reward scaling and realism audits (Nie and Li, 2024). Large-scale human-in-the-loop testing remains cost-prohibitive, averaging \$0.42 per labeled explanation (Tschiatschek et al., 2024).
Privacy, Compliance, and Safety Risks – Generative models leak PII through memorization or adversarial injection (Chizari, 2025; Smith et al., 2024). Differentially private fine-tuning (Li et al., 2024c; Bu et al., 2024), federated prompt learning (Tran et al., 2025; Hou et al., 2025), and secure retrieval-augmented generation (Koga et al., 2025) lower exposure but add 10–20 % latency. SMPC-backed inference enclaves and SGX isolated prompts narrow leakage channels yet limit model size (Johnson and Dey, 2024). New extraction attacks targeting personalization prompts achieve 18 % success even under DP noise (Rahman and Choi, 2025). DMA “right-to-explanation” mandates add extra logging and user-facing redaction complexity (Zou and Zhang, 2022).
Debugging and Observability Deficits – LLM pipelines lack token-level telemetry; silent regressions can persist until surfaced by downstream KPIs. Prompt lineage, streaming log replay, and token-attribution saliency maps are emerging, yet add 2 GB/day logs per million requests (Luo and Specia, 2024). Black-box causal probes help trace ranking shifts but require expensive interventional traffic (Schnabel et al., 2016). Safety dashboards incorporating toxic-span detection and PII scanners catch 92 % violations offline, but only 71 % in live traffic due to paraphrase drift (Lee and Zhou, 2025). Debugging culture and tooling lags behind deterministic retrievers.
Architectural Integration Overhead – Production recommenders rely on vector search and feature-based scoring. Plugging in LLMs requires async rerankers, RAG stacks, prompt stores, and rollback-safe deployment paths, each adding failure modes (Wu et al., 2024b). Multi-tenant isolation, memory pressure, and GPU scheduling remain open issues, especially in edge and on-device scenarios (Kim and Agarwal, 2025; Kang and Lee, 2024). Cross-service dependency graphs grow exponentially, complicating SRE ownership boundaries and raising MTTR by 1.4× (Nagy et al., 2025). End-to-end benchmarks for hybrid (LLM + vector) architectures are still nascent (Song and Bala, 2024).
Tackling these intertwined limitations—spanning latency, cost, hallucination, prompt robustness, drift, evaluation, privacy, observability, and architecture—is essential before LLM-enhanced recommenders can be deployed safely and economically at web scale.
6. Conclusion
LLMs represent a transformative shift in the architecture and design philosophy of recommender systems. By enabling contextual reasoning, zero-shot personalization, and multimodal representation learning, LLMs transcend the limitations of traditional matrix factorization, collaborative filtering, and feed-forward neural architectures. Rather than relying solely on historical interaction data and engineered features, LLM-augmented pipelines can synthesize user intent, simulate plausible behaviors, and generate semantically enriched item representations from natural language descriptions, reviews, and metadata. This capability significantly enhances performance in cold-start, long-tail, and rapidly evolving user contexts.
Throughout this paper, we examined how LLMs can be strategically integrated into various stages of industrial recommender pipelines—from candidate generation and reranking to feedback interpretation and synthetic evaluation. Our analysis demonstrated that LLMs not only offer improvements in semantic fidelity and explainability but also introduce opportunities for conversational and intent-driven recommendation paradigms. Techniques such as retrieval-augmented generation, prompt-based reranking, and hybrid collaborative-semantic scoring models exemplify how LLMs can coexist with traditional recommenders to achieve both robustness and expressivity.
However, this integration is not without its challenges. LLM-based systems pose significant barriers in latency-sensitive environments, where autoregressive decoding and token-based computation are at odds with sub-second serving constraints. Furthermore, issues such as prompt brittleness, hallucinated outputs, representation drift across model versions, and inconsistent personalization responses necessitate new debugging, logging, and evaluation toolchains. These technical hurdles are further complicated by operational constraints, including privacy compliance (e.g., GDPR, CCPA), prompt injection vulnerabilities, and the lack of interpretability guarantees in generative outputs. To realize the full potential of LLMs in recommender systems, future research must address the foundational tension between expressivity and efficiency. Robust prompt design and caching strategies, low-latency inference via distilled or quantized models, and standardized protocols for prompt versioning and auditability will be critical for production readiness. In parallel, new forms of evaluation that combine qualitative user modeling, counterfactual simulation, and LLM-driven assessments are needed to measure real-world effectiveness beyond traditional top-$k$ metrics.

### v2

@@
 Our work provides a foundation for researchers and engineers to reimagine recommender systems in the era of large language modeling, and highlights promising directions for future innovation.
+Figure 1.
+Taxonomy of recommender system challenges and corresponding LLM-based solutions.
+The framework categorizes challenges into four major areas—data-centric, modeling and algorithmic, evaluation, and privacy/security—each with sub-problems and LLM-driven strategies for addressing them.’=
 2. Evolution of Architectures in Recommender Systems
-The architectural landscape of recommender systems has undergone a remarkable transformation over the past two decades.
-Beginning with simple heuristic methods and collaborative filtering techniques, the field has progressively embraced more sophisticated paradigms powered by deep learning, graph neural networks, and, more recently, LLMs.
-This evolution reflects the growing demand for personalization, scalability, and contextual awareness in real-world applications.
-This section presents a chronological overview of key architectural advancements in recommender systems, highlighting the motivations, technical foundations, and practical implications of each generation.
-By tracing this progression, we aim to shed light on how the field has adapted to increasingly complex data modalities, user behaviors, and computational constraints.
-2.1. Traditional Approaches
-Heuristic-Based: Heuristic-based recommender systems represent some of the earliest strategies for generating item suggestions.
-These approaches rely on deterministic rules and statistical aggregations derived from historical data, rather than modeling latent user-item interactions.
-They are often used in large-scale systems for bootstrapping recommendations or serving as baseline models (Schafer et al., 2001; Bobadilla et al., 2018).
-A common heuristic is popularity-based recommendation, which ranks items based on global interaction frequency.
-For instance, the recommendation score $s_{i}$ for an item $i$ may be computed as:
-$s_{i}=\text{count}(i)=\sum_{u\in\mathcal{U}}\mathbb{I}(i\in I_{u})$
-where $\mathbb{I}(\cdot)$ is the indicator function, $I_{u}$ is the set of items interacted with by user $u$, and $\mathcal{U}$ is the set of all users.
-This formulation favors frequently interacted items, such as most-viewed videos or best-selling products.
-Another widely used rule is co-occurrence-based recommendation, which suggests items based on observed pairwise co-interactions.
-For example, an associative rule like “Users who bought or viewed item $i$ also bought or viewed item $j$” is operationalized using:
-$\text{sim}(i,j)=\frac{\text{count}(i,j)}{\sqrt{\text{count}(i)\cdot\text{count}(j)}}$
-where $\text{count}(i,j)$ is the number of users who interacted with both items $i$ and $j$, serving as a symmetric co-occurrence similarity metric (Sarwar et al., 2001).
-Despite their simplicity and scalability, heuristic methods are inherently non-personalized—they assume user homogeneity and do not adapt to individual preferences.
-As such, they suffer from popularity bias and lack context-awareness, making them less suitable for tasks requiring fine-grained personalization (Adomavicius and Tuzhilin, 2005).
-Nonetheless, their low computational cost and interpretability make them useful in cold-start scenarios where user history is sparse, in bootstrapping candidate sets in multi-stage recommender pipelines, and in serving fallback recommendations when more advanced model outputs are unreliable.
-In many production systems, heuristic scores are combined with learned models in later ranking stages or incorporated into ensemble frameworks for robustness.
-Figure 1.
-(i) Content-Based Recommendation: Recommends movies based on content features such as genre and cast similarity with previously watched movies.
-(ii) User-Based Collaborative Filtering: Recommends movies by identifying users with similar tastes and suggesting items liked by similar users.
-(Mallick, 2023)
-Collaborative Filtering: Collaborative Filtering (CF)  (Zhang et al., 2014) is a core methodology in recommender systems that predicts a user’s interest in an item by leveraging patterns of user-item interactions.
-The central assumption behind CF is the homophily principle: users who have shown similar interaction patterns in the past are likely to share future preferences.
-Unlike content-based methods, CF requires no side information or item/user metadata—it exclusively relies on interaction history, such as ratings, clicks, or purchases.
-Formally, given a sparse interaction matrix $\mathbf{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$ where $\mathcal{U}$ is the set of users and $\mathcal{I}$ is the set of items, the goal is to estimate the unobserved ratings or affinities $\hat{r}_{u,i}$ for a user-item pair $(u,i)$.
-There are two major paradigms of CF.
-User-based CF computes the similarity between the target user $u$ and other users $v\in\mathcal{U}$, and then aggregates the preferences of the $k$ most similar users $\mathcal{N}_{k}(u)$ to predict a score using the formula:
-$\hat{r}_{u,i}=\frac{\sum_{v\in\mathcal{N}_{k}(u)}\text{sim}(u,v)\cdot r_{v,i}}{\sum_{v\in\mathcal{N}_{k}(u)}|\text{sim}(u,v)|}$
-where $\text{sim}(u,v)$ may be computed using cosine similarity, Pearson correlation, or Jaccard similarity over the vectors $\mathbf{r}_{u}$ and $\mathbf{r}_{v}$.
-In contrast, item-based CF finds items similar to those the user has interacted with, using item-item similarities. The estimated score is:
-$\hat{r}_{u,i}=\frac{\sum_{j\in\mathcal{I}_{u}}\text{sim}(i,j)\cdot r_{u,j}}{\sum_{j\in\mathcal{I}_{u}}|\text{sim}(i,j)|}$
-where $\mathcal{I}_{u}$ is the set of items interacted with by user $u$, and similarity can again be cosine or adjusted cosine similarity.
-Common similarity functions include cosine similarity:
-$\text{sim}_{\text{cosine}}(i,j)=\frac{\mathbf{r}_{*,i}\cdot\mathbf{r}_{*,j}}{\|\mathbf{r}_{*,i}\|\cdot\|\mathbf{r}_{*,j}\|}$
-and Pearson correlation:
-$\text{sim}_{\text{Pearson}}(i,j)=\frac{\sum_{u}(r_{u,i}-\bar{r}_{i})(r_{u,j}-\bar{r}_{j})}{\sqrt{\sum_{u}(r_{u,i}-\bar{r}_{i})^{2}}\sqrt{\sum_{u}(r_{u,j}-\bar{r}_{j})^{2}}}$
-Despite its simplicity and interpretability, CF faces several challenges.
-Chief among them are data sparsity and the cold-start problem—both user and item-level—which hinder similarity computation.
-Real-time similarity computation can also become computationally expensive as the number of users or items grows.
-To mitigate these limitations, CF is often combined with content-based models in hybrid architectures.
-Furthermore, recent advances have proposed learning-based CF methods such as matrix factorization and neural embeddings.
-For instance, LightGCN (He et al., 2020) simplifies graph-based collaborative filtering by removing feature transformations and non-linearities, using only weighted summation of neighborhood embeddings:
-$\mathbf{e}_{u}^{(k)}=\sum_{i\in\mathcal{N}(u)}\frac{1}{\sqrt{|\mathcal{N}(u)||\mathcal{N}(i)|}}\cdot\mathbf{e}_{i}^{(k-1)}$
-This iterative propagation enhances higher-order connectivity, improving recommendation accuracy while maintaining linear complexity.
-Collaborative Filtering remains a fundamental building block in many industrial pipelines and serves as a baseline for benchmarking modern deep learning recommender systems (Li et al., 2023d; Rendle et al., 2020).
-Matrix Factorization: Matrix Factorization (MF)  (Zhang, 2022) is a class of latent factor models that projects users and items into a shared low-dimensional vector space to model user-item interactions.
-Given a user–item rating matrix $R\in\mathbb{R}^{m\times n}$, where $m$ is the number of users and $n$ is the number of items, the goal is to approximate $R$ by the product of two low-rank matrices:
-$R\approx\hat{R}=UV^{\top},$
-where $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$ are the user and item latent factor matrices, and $k\ll\min(m,n)$ is the number of latent dimensions.
-The optimization objective is typically to minimize the squared error over observed entries:
-$\min_{U,V}\sum_{(u,i)\in\mathcal{K}}(r_{u,i}-\mathbf{u}_{u}^{\top}\mathbf{v}_{i})^{2}+\lambda\left(\|\mathbf{u}_{u}\|^{2}+\|\mathbf{v}_{i}\|^{2}\right),$
-where $\mathcal{K}$ is the set of observed ratings, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent vectors for user $u$ and item $i$, respectively, and $\lambda$ is a regularization hyperparameter to prevent overfitting.
-Several variations of matrix factorization have been proposed to enhance recommendation accuracy and scalability.
-Singular Value Decomposition (SVD)  (Klema and Laub, 1980) applies truncated decomposition to the user-item rating matrix, projecting it into lower-dimensional latent factors; however, it often requires prior imputation of missing entries, which can introduce bias.
-Alternating Least Squares (ALS) offers a more scalable alternative by optimizing the matrix factorization objective through iterative updates: fixing one of the user or item matrices while solving a regularized least squares problem for the other (Zhou et al., 2008).
-This alternating procedure makes ALS particularly well-suited for large-scale and distributed environments, such as those in industrial settings.
-An extension to these approaches, SVD++, incorporates implicit feedback—such as clicks or views—alongside explicit ratings to refine user representations.
-The predicted rating in SVD++ is given by:
-$\hat{r}_{u,i}=\mu+b_{u}+b_{i}+\left(\mathbf{u}_{u}+\sum_{j\in\mathcal{I}_{u}}\frac{y_{j}}{\sqrt{|\mathcal{I}_{u}|}}\right)^{\top}\mathbf{v}_{i},$
-where $\mu$ is the global average rating, $b_{u}$ and $b_{i}$ are user and item-specific biases, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent factors for user $u$ and item $i$, respectively, and $y_{j}$ represents the embedding of item $j$ contributing to user $u$’s implicit feedback profile.
-Matrix factorization has become the backbone of many collaborative filtering systems due to its balance between expressiveness and scalability.
-Its popularity surged during the Netflix Prize competition, and it continues to serve as a robust baseline for collaborative filtering tasks (Koren et al., 2009).
-2.2. Content-Based and Hybrid Systems
-Content-Based Filtering: Content-based filtering (CBF)  (Wang et al., 2024d) recommends items to a user by analyzing the similarity between the items’ attributes and the user’s historical preferences (Lops et al., 2011).
-Unlike collaborative methods, CBF operates independently of other users and relies solely on item metadata—such as genre, tags, textual descriptions, or visual features.
-A user profile $\mathbf{u}$ is typically constructed by aggregating the feature vectors of previously liked or interacted items.
-For a given item $i$ with a feature vector $\mathbf{x}_{i}$, the predicted preference score is computed using a similarity function:
-$\hat{r}_{u,i}=\text{sim}(\mathbf{u},\mathbf{x}_{i})$
-Common choices for sim include cosine similarity:
-$\text{sim}_{\text{cosine}}(\mathbf{u},\mathbf{x}_{i})=\frac{\mathbf{u}^{\top}\mathbf{x}_{i}}{\|\mathbf{u}\|\cdot\|\mathbf{x}_{i}\|}$
-In text-based domains, TF-IDF representations or pre-trained language model embeddings are often used to define $\mathbf{x}_{i}$, while in visual domains, image embeddings from CNNs or vision transformers are leveraged (Mouhiha et al., 2024).
-While CBF models are robust to cold-start users (as long as item features are known), they struggle with over-specialization and limited discovery of novel content.
-Hybrid Models: Hybrid recommender systems combine content-based and collaborative filtering paradigms to overcome the limitations of either approach in isolation (Burke, 2002).
-The fusion can occur at various levels—model, feature, or score—and often improves robustness, especially in sparse or cold-start scenarios.
-LightFM (Kula, 2015) incorporates both collaborative signals (user-item interactions) and content features (metadata) into a unified latent factor model.
-Its objective is to learn user and item embeddings from both sources:
-$\hat{r}_{u,i}=\mathbf{u}_{u}^{\top}\mathbf{v}_{i}+\mathbf{u}_{u}^{\top}\mathbf{f}_{i}+\mathbf{f}_{u}^{\top}\mathbf{v}_{i},$
-where $\mathbf{f}_{u}$ and $\mathbf{f}_{i}$ are feature representations of user and item metadata respectively.
-DeepFM (Guo et al., 2017) further extends this by combining a Factorization Machine (FM) layer for low-order feature interactions with a deep neural network for high-order interactions.
-while the deep component captures non-linear patterns across raw features via multi-layer perceptrons.
-These hybrid architectures provide stronger generalization across heterogeneous data sources and are widely used in large-scale recommendation engines.
-2.3. Deep Learning Architectures
-Neural Collaborative Filtering (NCF): Neural Collaborative Filtering (NCF)  (He et al., 2017a) is a deep learning-based extension of matrix factorization that captures complex, nonlinear user–item interaction patterns.
-Instead of modeling the dot product between latent user and item vectors, NCF uses a Multi-Layer Perceptron (MLP) to learn a nonlinear interaction function:
-$\hat{r}_{u,i}=\phi(\mathbf{u}_{u},\mathbf{v}_{i})=\text{MLP}([\mathbf{u}_{u};\mathbf{v}_{i}]),$
-where $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the embedding vectors for user $u$ and item $i$, respectively, and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation.
-The architecture enables the model to approximate arbitrary functions between user and item representations.
-The NeuMF model (He et al., 2017b) combines Generalized Matrix Factorization (GMF) and MLP-based modeling via a fusion layer to benefit from both linear and nonlinear signals.
-Figure 2.
-Neural Collaborative Filtering (NCF) architecture: One-hot user and item inputs are embedded into latent vectors, passed through neural layers, and produce a predicted score $\hat{y}_{ui}$ trained against the target $y_{ui}$ (He et al., 2017a)
-Wide & Deep Models: The Wide & Deep architecture (Cheng et al., 2016) aims to capture both low- and high-order feature interactions by combining a linear model (”wide”) and a deep neural network (”deep”) within a joint framework:
-$\hat{y}=\sigma(\mathbf{w}^{\top}\mathbf{x}+\mathbf{a}^{\top}f_{\text{deep}}(\mathbf{x})),$
-where $\mathbf{x}$ is the feature vector, $\mathbf{w}$ are the wide (memorization) weights, and $f_{\text{deep}}(\mathbf{x})$ represents the output of the deep network.
-This hybrid design enables the model to both remember frequent co-occurrences (e.g., user clicked item A and B) and generalize to unseen feature combinations.
-It was deployed in production at Google Play for app recommendations, demonstrating substantial improvements in personalization and click-through rates.
-2.4. Sequence-Aware Models
-As user behavior in many applications is inherently sequential (e.g., streaming, shopping, or reading), capturing temporal dependencies becomes crucial.
-Sequence-aware recommender systems model the order and dynamics of user-item interactions over time, enabling next-item prediction based on behavioral context.
-RNN-Based Models: Recurrent Neural Networks (RNNs)  (Schmidt, 2019), particularly Gated Recurrent Units (GRUs) and Long Short-Term Memory (LSTM) networks, have been used to model user behavior sequences.
-GRU4Rec (Hidasi et al., 2016) was one of the earliest models to apply RNNs to session-based recommendation, treating the problem as a sequence prediction task.
-Let $x_{t}$ be the embedding of the item interacted with at time $t$. The GRU updates its hidden state $h_{t}$ using:
-$h_{t}=\text{GRU}(x_{t},h_{t-1}),$
-and predicts the next item via:
-$\hat{y}_{t+1}=\text{softmax}(Wh_{t}+b),$
-where $W$ and $b$ are learnable parameters.
-GRU4Rec enables learning from both short-term (within-session) and long-term user preferences and has shown superior performance in session-based scenarios compared to traditional CF methods.
-Transformer-Based Models: Transformer-based  (Vaswani et al., 2023) models replace the recurrence in RNNs with self-attention mechanisms, allowing better modeling of long-range dependencies and parallelization.
-SASRec (Kang and McAuley, 2018) applies a unidirectional self-attention encoder to model sequential interactions:
-$\mathbf{H}=\text{SelfAttention}(\mathbf{X}+\mathbf{P}),$
-where $\mathbf{X}$ is the sequence of item embeddings and $\mathbf{P}$ is the positional encoding. The model predicts the next item using the last hidden state:
-$\hat{y}_{t+1}=\text{softmax}(Wh_{t}).$
-BERT4Rec (Sun et al., 2019) introduces bidirectional attention with a masked item prediction objective, inspired by the BERT architecture in NLP.
-This enables the model to learn contextual dependencies from both past and future items in a sequence.
-TiSASRec (Li et al., 2020) extends SASRec by incorporating both absolute and relative time intervals into attention weights, enhancing temporal modeling in user behavior:
-$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}+T_{\text{abs}}+T_{\text{rel}}}{\sqrt{d_{k}}}\right)V,$
-where $T_{\text{abs}}$ and $T_{\text{rel}}$ encode absolute and relative temporal distances, respectively.
-Overall, transformer-based models have surpassed RNNs in many sequence-aware recommendation tasks due to their ability to model long-term dependencies and leverage parallel computation.
-2.5. Graph-Based Recommenders
-Graph-based recommender systems  (Wang et al., 2021a) represent user–item interactions as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathcal{E})$, where $\mathcal{U}$ is the set of users, $\mathcal{I}$ is the set of items, and edges $\mathcal{E}$ correspond to observed interactions (e.g., clicks, ratings, purchases).
-This graph structure enables the modeling of high-order connectivity and collaborative signals beyond immediate neighbors, capturing richer semantics than traditional matrix factorization.
-Figure 3.
-Illustration of a heterogeneous graph-based recommender system integrating three components: the Social Relation Graph (user-user links), the User-Item Interaction Graph (explicit and implicit feedback such as clicks and ratings), and the Domain Knowledge Graph (item attributes like OS and color).
-This structure facilitates richer context-aware and socially-informed recommendations (Kipf and Welling, 2017)
-Graph Convolutional Networks (GCNs) have been widely adopted for such modeling  (Kipf and Welling, 2017).
-A GCN propagates information between nodes through message passing, updating the representation of a node by aggregating features from its neighbors.
-In a simplified GCN layer:
-$\mathbf{h}_{v}^{(k)}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{W}^{(k)}\mathbf{h}_{u}^{(k-1)}\right),$
-where $\mathbf{h}_{v}^{(k)}$ is the embedding of node $v$ at the $k^{\text{th}}$ layer, $\mathbf{W}^{(k)}$ is a learnable weight matrix, and $\mathcal{N}(v)$ is the set of neighbors of $v$.
-NGCF (Neural Graph Collaborative Filtering) (Ying et al., 2018) builds on this idea by combining collaborative filtering with graph convolution.
-It explicitly models user–item interactions via non-linear message propagation.
-At each layer, it computes:
-$\mathbf{e}_{v}^{(k)}=\sigma\left(\mathbf{W}_{1}^{(k)}\mathbf{e}_{v}^{(k-1)}+\mathbf{W}_{2}^{(k)}\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}\right),$
-where $\mathbf{e}_{v}^{(k)}$ is the embedding of user or item $v$ at layer $k$.
-LightGCN (He et al., 2020) simplifies NGCF by removing the nonlinear activation and transformation weights, arguing that they may over-complicate the propagation. LightGCN computes embeddings as:
-$\mathbf{e}_{v}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{v}^{(k)},\quad\text{where}\quad\mathbf{e}_{v}^{(k)}=\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}.$
-PinSage (Ying et al., 2018) extends this concept for large-scale web-scale recommendation systems (e.g., Pinterest).
-It combines random walks with graph convolutions and incorporates node features using neighborhood sampling and importance-based aggregation.
-Graph-based models are especially effective at capturing high-order collaborative signals and performing inductive reasoning on unseen nodes.
-They have become foundational in modern recommenders deployed in large-scale platforms due to their flexibility and scalability.
-2.6. Large-Scale Retrieval Models
-Large-scale retrieval models are designed to efficiently handle recommendation tasks at industrial scale, where the corpus of items may reach millions or billions.
-These models prioritize fast candidate generation and scalability while maintaining reasonable recommendation quality.
-Figure 4.
-Hybrid ranking architecture combining two-tower and pairwise models.
-User-specific item representations and shallow embeddings are processed through fusion scoring to generate the final ranked list.
-(Yuan et al., 2024).
-Two-Tower Models: Two-tower architectures, also known as dual-encoder models  (Yuan et al., 2024), are among the most commonly deployed retrieval methods in production systems.
-These models learn independent embedding functions for users and items:
-$\mathbf{e}_{u}=f_{\theta}(u),\quad\mathbf{e}_{i}=g_{\phi}(i)$
-where $f_{\theta}$ and $g_{\phi}$ are neural networks parameterized by $\theta$ and $\phi$, respectively.
-The final relevance score between user $u$ and item $i$ is computed as the inner product or cosine similarity:
-$s(u,i)=\langle\mathbf{e}_{u},\mathbf{e}_{i}\rangle$
-This separation allows precomputing item embeddings and indexing them via Approximate Nearest Neighbor (ANN) techniques, enabling sublinear retrieval times.
-The YouTube DNN model (Covington et al., 2016) is an influential example of a two-tower retrieval system.
-It uses multi-layer perceptrons (MLPs) for both user and item towers, trained with sampled softmax to optimize for the target watch history.
-Another widely used architecture is Facebook’s Deep Learning Recommendation Model (DLRM) (Naumov et al., 2019), which combines dense and sparse features from user-item pairs and leverages embedding lookups and interaction layers to compute the final relevance score.
-These models form the first stage in multi-stage ranking pipelines.
-After candidate generation via ANN search over user-item embeddings, a more expressive model (e.g., gradient boosting, transformers) typically performs re-ranking over the retrieved shortlist.
-Due to their efficiency and deployment advantages, two-tower retrieval models are considered foundational in modern industrial recommender systems.
+Recommender systems have progressed from heuristic rules to collaborative filtering, latent factorization, neural and graph-based models, and finally industrial-scale retrieval frameworks.
+Each paradigm enhanced expressiveness, scalability, and the ability to capture complex dependencies.
+2.1. From Heuristics to Latent Factor Models
+Early recommenders used deterministic heuristics such as popularity- and co-occurrence-based rules (Schafer et al., 2001; Bobadilla et al., 2018; Sarwar et al., 2001), which scale well but lack personalization.
+Collaborative Filtering (CF) (Zhang et al., 2014) introduced personalization by exploiting similarity in interaction patterns between users or items.
+While simple and interpretable, CF suffers from sparsity and cold-start issues.
+Matrix Factorization (MF) (Klema and Laub, 1980; Koren et al., 2009) addressed scalability by embedding users and items into a shared low-dimensional space, with predictions computed via inner products of latent vectors.
+Variants such as ALS (Zhou et al., 2008) and SVD++ (Koren et al., 2009) incorporated implicit feedback and became widely adopted in industrial systems due to their balance of accuracy and efficiency.
+2.2. Hybrid and Neural Extensions
+Content-Based Filtering (CBF) (Lops et al., 2011) enriched recommendations with item attributes such as textual, categorical, or visual embeddings, making it more robust to cold-start users.
+However, CBF tends to overspecialize.
+Hybrid models (Burke, 2002) combined collaborative and content signals, exemplified by LightFM (Kula, 2015) and DeepFM (Guo et al., 2017), which integrate metadata with learned embeddings.
+Deep learning extended this line with Neural Collaborative Filtering (NCF) (He et al., 2017), which replaces linear interaction functions with nonlinear neural networks, and Wide & Deep (Cheng et al., 2016), which jointly optimizes memorization and generalization.
+These architectures established the basis for modern industrial recommenders.
+2.3. Sequential and Graph-Based Models
+User behavior often exhibits temporal and contextual dynamics.
+Sequence-aware recommenders such as GRU4Rec (Hidasi et al., 2016) applied recurrent neural networks to session data, while transformer-based approaches (e.g., SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSASRec (Li et al., 2020)) leveraged self-attention to capture long-range dependencies and temporal intervals.
+These models outperform traditional CF in sequential domains such as streaming and e-commerce.
+In parallel, graph-based recommenders (Wang et al., 2021a) represented interactions as bipartite graphs.
+Neural Graph Collaborative Filtering (NGCF) (Ying et al., 2018) propagated embeddings through nonlinear message passing, while LightGCN (He et al., 2020) simplified this to linear aggregation, improving both accuracy and scalability.
+PinSage (Ying et al., 2018) scaled graph-based recommendation to billions of nodes using random walks and sampling, enabling deployment in web-scale platforms like Pinterest.
+2.4. Industrial-Scale Retrieval Frameworks
+With item catalogs often exceeding billions, retrieval efficiency became a dominant concern.
+Two-tower architectures (Yuan et al., 2024) learn independent embeddings for users and items, with relevance computed as the inner product between vectors.
+Pre-computed item embeddings can be indexed using Approximate Nearest Neighbor (ANN) search, enabling sublinear retrieval times.
+The YouTube DNN (Covington et al., 2016) demonstrated the effectiveness of deep two-tower retrieval with sampled softmax, while Facebook’s DLRM (Naumov et al., 2019) integrated sparse and dense features with cross-feature interaction layers.
+These retrieval modules form the candidate generation stage in multi-stage pipelines, with re-ranking typically handled by more expressive models such as gradient boosting or transformers.
 3. Key Components of Recommender System Pipelines
@@
 Cross-Domain Transfer
-(Zhang et al., 2021a; Li et al., 2023a; Zhai, 2024; Xu et al., 2023b; Kang et al., 2023; Dai et al., 2023)
+(Zhang et al., 2021a; Li et al., 2023a; Zhai, 2024; Xu et al., 2023b; Kang et al., 2023; Dai et al., 2023b)
 (Huang and Huang, 2024; Gao et al., 2021; Gomez-Uribe and Hunt, 2021; Asai and Hajishirzi, 2020)
@@
 (Ranasinghe et al., 2025; Alayrac et al., 2022; Tian et al., 2025; Yu et al., 2022)
-(Wang et al., 2024c; Finn et al., 2017; Lee et al., 2019)
+(Wang et al., 2024c; Finn et al., 2017a; Lee et al., 2019)
 (Vajjala et al., 2024; Yi et al., 2023; Liu et al., 2025b; Vats et al., 2024)
@@
 (Lu et al., 2022, 2023)
-(Sapkota et al., 2025; Zhang et al., 2025d)
+(Sapkota et al., 2025; Zhang et al., 2025e)
 (Radford et al., 2021; Zhou et al., 2025; Chen et al., 2022)
@@
 (Zhang et al., 2021a; Kang et al., 2023)
-(Vasile et al., 2022; Dai et al., 2023; Wu et al., 2024a)
+(Vasile et al., 2022; Dai et al., 2023b; Wu et al., 2024a)
 (Chung et al., 2022; Wang et al., 2023a)
@@
 Language-Guided Reweighting
-(Dai et al., 2023; Zhai, 2024; Zhang et al., 2021a)
+(Dai et al., 2023b; Zhai, 2024; Zhang et al., 2021a)
 (Wong and Tanaka, 2025)
@@
 The sparsity of user–item signals and the lack of initialization paths for unseen entities make cold start a persistent bottleneck in the design of effective recommender systems (Schein et al., 2002; Park and Tuzhilin, 2006).
-LLMs offer a paradigm shift in mitigating the cold start problem by leveraging their ability to process rich textual and contextual information without requiring historical interaction data (Dai et al., 2023; Wang et al., 2024b).
+LLMs offer a paradigm shift in mitigating the cold start problem by leveraging their ability to process rich textual and contextual information without requiring historical interaction data (Dai et al., 2023b; Wang et al., 2024b).
 Unlike traditional collaborative filtering methods that depend on dense user–item matrices, LLMs can infer relevance through semantic reasoning, prompt-based generation, and cross-modal understanding.
@@
 Retrieval-Augmented Generation (RAG) – RAG-based methods enhance LLMs with a retrieval module that selects relevant items or documents based on the input query.
-In recommender systems, this architecture decouples memorization from inference, allowing the model to attend to similar items, behavioral patterns, or metadata even when the target entity is entirely new or sparsely observed (Kang et al., 2023; Dai et al., 2023).
+In recommender systems, this architecture decouples memorization from inference, allowing the model to attend to similar items, behavioral patterns, or metadata even when the target entity is entirely new or sparsely observed (Kang et al., 2023; Dai et al., 2023b).
 By combining parametric knowledge from the LLM with non-parametric access to a large external index, RAG frameworks enable more accurate, controllable, and interpretable recommendation outputs.
@@
 It enables interaction-rich scenarios even in the absence of dense interaction histories, facilitating few-shot generalization in cold-start settings  (Liu et al., 2024).
-Notably, recent systems like PromptRec (Dai et al., 2023), PROMO (Jiang et al., 2024), and OpenAI’s plugin-based RecSys prototype (Li et al., 2023a) have operationalized such strategies in both online and offline settings.
+Notably, recent systems like PromptRec (Dai et al., 2023b), PROMO (Jiang et al., 2024), and OpenAI’s plugin-based RecSys prototype (Li et al., 2023a) have operationalized such strategies in both online and offline settings.
 These systems dynamically compose prompts using structured user/item attributes and contextual cues (e.g., time, location, browsing intent), effectively merging retrieval and generation within a single architecture.
@@
 As LLMs continue to improve in understanding long-tail user intents and compositional queries, prompt-based conditioning is emerging as a powerful mechanism for building controllable, transparent, and context-aware recommender systems.
-Multimodal Embedding Synthesis – When textual information alone is insufficient for high–fidelity personalization, LLMs can be situated at the center of a multimodal pipeline that fuses visual, structured, and behavioral signals into a unified representation  (Ranasinghe et al., 2025).
+Multimodal Embedding Synthesis –When textual information alone is insufficient for high–fidelity personalization, LLMs can be situated at the center of a multimodal pipeline that fuses visual, structured, and temporal signals  Figure 3).
+d behavioral signals into a unified representation  (Ranasinghe et al., 2025).
 A generic fusion template is:
@@
 where $x_{i}^{\text{text}}$ (e.g., product title, description), $x_{i}^{\text{img}}$ (primary and auxiliary images), and $x_{i}^{\text{meta}}$ (price, brand, category, timestamp, CTR priors) are embedded by modality-specific encoders and subsequently aligned in a shared latent space.
+Figure 3.
+Overview of the Proposed Multi-Modal Sequence Learning Framework.(a) The model processes timestamp, ID, text, and image sequences using BERT, ViT, and ID embeddings, followed by Interactive and Temporal MoE modules for expert routing.
+Outputs are fed into Transformers for prediction.
+(b) Multi-task learning includes category prediction, contrastive learning on ID embeddings, and placeholder-based contrastive learning across modalities  (Zhang et al., 2025b).
 State-of-the-art LLMs such as Flamingo (Alayrac et al., 2022) use cross-attention layers to inject image token embeddings into frozen language blocks, while lighter adapters or multi-head cross-modal transformers can accomplish a similar goal with lower inference cost.
@@
 Amazon’s Multimodal RecSys (Tian et al., 2025) employs hierarchical fusion: fine-tuned CLIP encoders supply image vectors, BERT-based encoders supply text, and catalog metadata is featurized via embedding tables.
-A shallow cross-modal transformer merges these signals and outputs $\mathbf{q}_{i}$ to a traditional ANN service, yielding a $\sim$5–10lift in hit-rate for products with <5 historical interactions  (Xu et al., 2023c).
+A shallow cross-modal transformer merges these signals and outputs $\mathbf{q}_{i}$ to a traditional ANN service, yielding a $\sim$5–10% lift in hit-rate for products with <5 historical interactions  (Xu et al., 2023c).
 Similar blueprints are emerging in streaming platforms that blend thumbnails, transcripts, and user watch-history to build rich item sketches for zero-shot retrieval  (Xu et al., 2022).
@@
 Meta-Learning for Fast Adaptation – Meta-learning provides a framework for enabling rapid personalization by training models to adapt quickly to new users or items using limited data  (Wang et al., 2024c).
-When combined with LLMs, meta-learning techniques such as Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) allow systems to fine-tune user- or item-specific representations via a small number of gradient updates.
+When combined with LLMs, meta-learning techniques such as Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017a) allow systems to fine-tune user- or item-specific representations via a small number of gradient updates(see Figure 4.
 For a new user task $\mathcal{T}_{u}$, the model parameters are adapted as follows:
@@
 These methods also support continual learning across user cohorts, enabling fast model updates without overfitting or catastrophic forgetting, and are particularly well-suited for dynamic environments like e-commerce or content streaming platforms  (Hamedi et al., 2025).
+Figure 4.
+The image compares MAML-based few-shot adaptation (left) to standard fine-tuning without MAML (right) for a sine wave regression task.
+MAML generalizes well, even in regions without data, by capturing the underlying periodic structure.
+In contrast, the pretrained model struggles to adapt from few examples, failing to extrapolate due to conflicting gradients from pretraining (Finn et al., 2017b)
 Cross-Domain Transfer via Language – LLMs pretrained on diverse and heterogeneous corpora can facilitate cross-domain recommendation by leveraging their semantic understanding to bridge disparate user-item interaction spaces  (Vajjala et al., 2024; Yi et al., 2023; Liu et al., 2025b; Vats et al., 2024).
@@
 Recent approaches such as PromptRec and LLMRec show that this strategy improves recommendation relevance in cross-modal and multilingual setups.
-Additionally, systems like CrossAligner (Gritta et al., 2022) use contrastive fine-tuning to align embeddings across domains, while  (Zhang et al., 2025c) demonstrate strong performance on cold-start item transfer using instruction-following LLMs.
+Additionally, systems like CrossAligner (Gritta et al., 2022) use contrastive fine-tuning to align embeddings across domains, while  (Zhang et al., 2025d) demonstrate strong performance on cold-start item transfer using instruction-following LLMs.
 This line of work opens avenues for applying pretraining-era language representations to cross-domain personalization at scale, eliminating the need for costly alignment mappings or dual-domain retraining.
@@
 (Liu et al., 2025c)
-(a) (i)
-(b) (ii)
-Figure 7.
-(i) LLM-Based Semantic Matching.
-The LLM converts textual metadata of new users and new items into dense embeddings.
-These embeddings are semantically matched via dot product to generate recommendation scores, enabling cold-start recommendations without historical interaction data.
-(ii) LLM-Based Personalization Techniques.
-The figure highlights four core strategies used by Large Language Models to enhance personalization: (1) Instruction-Tuned Generalization ensures global instruction-following; (2) Prompt-Tuned Personalization leverages embedded user feedback; (3) Behavioral Diversity is enabled through generative modeling; and (4) Multitask Prompting supports learning from multiple objectives.
-Together, these approaches enable more adaptive and user-aligned recommendation systems.
 4.1.2. Data Sparsity
@@
 This soft modality selection is realized implicitly, without explicit gating or masking heuristics  (Cai et al., 2025).
-Unified Representation through Instruction-Tuned LLMs – Given a structured prompt $\texttt{Prompt}_{i}$ that serializes all modality features (e.g., title, description, image caption, metadata, and recent user feedback), an instruction-tuned LLM can embed this multimodal input using a shared semantic space.
+Figure 7.
+An instruction-tuned LLM framework that unifies visual and textual inputs into a shared semantic space.
+The image encoder extracts visual embeddings, which are fused with textual instructions via a Q-Former module using self-attention and cross-attention.
+The combined representation is processed by the LLM to generate a response, enabling cross-modal reasoning without separate modality-specific towers.
+(Dai et al., 2023a)
+Unified Representation through Instruction-Tuned LLMs – Given a structured prompt $\texttt{Prompt}_{i}$ that serializes all modality features (e.g., title, description, image caption, metadata, and recent user feedback), an instruction-tuned LLM(see Figure 7 can embed this multimodal input using a shared semantic space.
 The dense item representation is computed as:
@@
 Cross-Modal Alignment via Pretraining – Large-scale pretraining on multimodal corpora (e.g., alt-text–image pairs, metadata–review co-occurrences) enables LLMs to implicitly align semantically related content across modalities  (Sapkota et al., 2025).
-These pretrained models induce a joint latent space where concepts are consistently grounded across different input forms  (Zhang et al., 2025d).
+These pretrained models induce a joint latent space where concepts are consistently grounded across different input forms  (Zhang et al., 2025e).
 At inference, when certain modalities are unavailable, the model can still generate meaningful representations by conditioning on the observed subset $x_{i}^{\text{obs}}\subset x_{i}$:
@@
 The model fills gaps by relying on shared linguistic anchors (e.g., a product’s visual properties inferred from its textual reviews or structured tags).
-This alignment strategy originated in vision-language models like CLIP  (Radford et al., 2021) and ALIGN  (Jia et al., 2021), and has been adapted for recommendation in PaLI (Chen et al., 2022), CoCa-2Rec (Wang et al., 2024e), and OmniFM (Zhou et al., 2025), where the model generalizes across image-text-review-metadata domains with minimal supervision.
+This alignment strategy originated in vision-language models like CLIP  (Radford et al., 2021) and ALIGN  (Jia et al., 2021), and has been adapted for recommendation in PaLI (Chen et al., 2022), CoCa-2Rec (Wang et al., 2024d), and OmniFM (Zhou et al., 2025), where the model generalizes across image-text-review-metadata domains with minimal supervision.
 Multimodal Imputation via Generative Inference – In sparse environments, generative LLMs can hallucinate plausible approximations for missing modalities based on contextual conditioning.
@@
 First, a lightweight candidate generator $g_{\text{retrieval}}$ retrieves a small set $\mathcal{C}(u)$ from the item pool, followed by a heavier ranking model $f_{\text{rank}}$:
-[…]
+(3) $\text{Top-}k=\arg\max_{i\in\mathcal{C}(u)}f_{\text{rank}}(u,i).$
 While this reduces overall latency, the challenge persists when incorporating deep user or item encoders. Let’s see how scalability challenge can be addressed using LLMs.
@@
 While traditional DNNs can be distilled, LLM-based distillation uniquely preserves instruction-following, multi-domain reasoning, and natural language grounding.
-Moreover, LLMs support multi-task distillation (e.g., joint training on ranking, generation, and explanation), enabling compressed models to generalize across different RecSys objectives (Dai et al., 2023; Liu et al., 2023c).
+Moreover, LLMs support multi-task distillation (e.g., joint training on ranking, generation, and explanation), enabling compressed models to generalize across different RecSys objectives (Dai et al., 2023b; Liu et al., 2023c).
 Prompt-Efficient Inference – Instead of full fine-tuning, prompting leverages frozen LLMs through task-specific prompts, reducing training and inference costs.
@@
 The hybrid setup enables real-time ranking with rich personalization, especially useful for high-value sessions where semantic intent understanding is critical.
-Sparse Activation Architectures – Sparse MoE (Mixture of Experts) models activate only a subset of parameters during inference, scaling to billions of weights without proportional cost (Lepikhin et al., 2021; Zoph et al., 2022).
+Figure 8.
+Illustration of a Sparse Mixture-of-Experts (MoE) layer where a gating network dynamically routes input tokens to a subset of experts.
+Only the top-k experts are activated per input, enabling efficient computation and scalability within Transformer architectures.
+(Shazeer et al., 2017)
+Sparse Activation Architectures – Sparse Mixture of Experts(see Figure 8) models activate only a subset of parameters during inference, scaling to billions of weights without proportional cost (Lepikhin et al., 2021; Zoph et al., 2022).
 For recommender systems, LLM-based MoE allows selective activation based on user intent or domain — as shown in GLaM (Du et al., 2022), Switch-LLM (Fedus et al., 2022), and V-MoE (Riquelme et al., 2021), which uses sparsity to optimize vision transformer inference.
@@
 A small $\lambda$ skews the model towards head items, suppressing cold and long-tail entities.
+The following subsections discuss how long tail modeling can be solved using LLMs.
 Content-Enriched Generation via LLMs – Instruction-tuned and generative LLMs can amplify long-tail items by transforming sparse metadata—such as item title, genre, and a few tags—into rich natural language content.
@@
 Retrieval-Augmented Tail Expansion – Tail items often suffer from cold-start issues due to minimal or no user engagement.
-Retrieval-Augmented Generation (RAG) frameworks alleviate this by fetching external content—such as similar item reviews, frequently co-viewed products, or question–answer forums—that offer surrogate interaction signals.
+Retrieval-Augmented Generation (RAG) frameworks alleviate this by fetching external content such as similar item reviews, frequently co-viewed products, or question–answer forums that offer surrogate interaction signals.
 These retrieved contexts act as priors, conditioning the LLM to produce recommendations or relevance scores with greater semantic grounding.
@@
 This paradigm has gained real-world traction in platforms like Amazon and Spotify (Wang et al., 2023b; Vasile et al., 2022), where hybrid systems interleave retrieval and generation in ranking.
-Public systems such as RAG (Lewis et al., 2021) and PromptRec (Dai et al., 2023) provide open-source implementations of this approach.
+Public systems such as RAG (Lewis et al., 2021) and PromptRec (Dai et al., 2023b) provide open-source implementations of this approach.
 RAG mechanisms also allow continual updates to item knowledge without re-training the LLM, offering scalability for dynamic catalogs.
@@
 Template-bound prompting offers improved compliance, while full LLM generation requires enhanced auditing and safety layers
-Figure 10.
-Illustration of Offline-Online Metric Divergence.
-This figure highlights the gap between offline evaluation—using historical logged data and metrics like NDCG, Recall@K, and MAP—and online deployment performance with real users, which is measured through business-critical KPIs such as click-through rate (CTR), average revenue per user (ARPU), session duration, and long-term user retention.
-This divergence presents a key challenge in validating recommender systems effectively.
 4.3.1. Offline-Online Gap
@@
 For each candidate $x_{i}$, a generative model produces a soft score:
-[…]
+$\tilde{r}_{ui}^{\text{LLM}}=\text{sigmoid}(\text{LLM}(x_{i},c_{u},\texttt{``How relevant is this item?''}))$
 These scores can be aggregated to construct hybrid metrics:
@@
 Generative Relevance Scoring How relevant is [item] for this user given their context?
-Behavior-Level Satisfaction The user clicked $x\_1$, skipped $x\_2$, then bounced. How satisfying was this session?
+Behavior-Level Satisfaction The user clicked x_1x\_1, skipped x_2x\_2, then bounced. How satisfying was this session?
 Agentic User Simulation You are a user interested in cooking. You are shown these recipes: [...]. What do you do next?
@@
 A representative prompt is shown in Table 6.
-This is especially valuable in semi-supervised learning regimes, where high-coverage but noisy labels can bridge the sparsity gap and enhance downstream recall (Dai et al., 2023; Zhai, 2024; Zhang et al., 2021a).
+This is especially valuable in semi-supervised learning regimes, where high-coverage but noisy labels can bridge the sparsity gap and enhance downstream recall (Dai et al., 2023b; Zhai, 2024; Zhang et al., 2021a).
 Instruction-Tuned Label Imputation - Instruction-following LLMs can reason about partial conversion traces and historical behavior patterns to impute likely outcomes.
@@
 User replies are encoded as
-[…]
+$\mathbf{u}_{\text{long}}=f_{\text{LLM}}(\texttt{user\_reply}),$
 and injected into evolving user models that steer ranking toward durable interests.
@@
 $\mathcal{J}_{\text{sim}}=\sum_{t=0}^{T}\gamma^{t}\,\hat{\mathcal{R}}_{t},$
-where each ^$\hat{\mathcal{R}}_{t}$ is predicted by the LLM under a hypothesized slate.
+where each $\hat{\mathcal{R}}_{t}$ is predicted by the LLM under a hypothesized slate.
 User-simulator frameworks such as SimRec (Chen et al., 2023b) and UserGPT-Sim (Zheng et al., 2023) report strong correlations (0.82–0.87) between simulated and real A/B outcomes.
@@
 4.4.4. Data Access and Governance
-Balancing innovation with data security requires robust data governance.
-Access control systems that prevent unauthorized access must coexist with tools that enable rapid development and experimentation.
-Companies like LinkedIn and Airbnb have built governance platforms that include role-based access control (RBAC), automated policy enforcement, and secure experimentation sandboxes (Liu et al., 2021).
-These systems help enforce compliance while supporting agility.
-Teams should adopt scalable governance frameworks that integrate access controls directly into ML and data infrastructure, ensuring accountability without obstructing productivity.
-Data Minimization via LLMs: One of the most promising applications of LLMs is their ability to reduce dependence on personally identifiable information (PII) by extracting meaningful signals from anonymized or high-level interaction data.
-Instead of collecting demographic features or granular histories, LLMs can infer user interests through textual representations of behavior, public reviews, or short intent prompts.
-This reduces the risk surface area while maintaining personalization effectiveness.
-The goal is to minimize the quantity of personal data needed for effective recommendation:
-$Q(\text{Rec})\approx Q_{\text{LLM}}(\text{Rec}\mid\text{anonymized data})$
-Here, $Q(\text{Rec})$ represents recommendation quality. If the quality from the LLM using anonymized inputs approximates that from full PII, the system has achieved successful data minimization.
-Recent work has shown that privacy-preserving techniques can be successfully applied to large language models in recommendation settings.
-One promising direction involves federated prompt learning with differential privacy, which can achieve performance comparable to centralized training while ensuring strong user-level protections  (Tran et al., 2025).
-Another approach leverages synthetic user simulations generated by LLMs, reducing reliance on sensitive behavioral logs.
-Additional strategies include token-level obfuscation to filter private information before inference, and private fine-tuning methods that mitigate membership inference risks without significantly compromising model accuracy  (Zhang et al., 2025b).
-Instructable Privacy Filters: LLMs can also be explicitly guided—via prompt templates or fine-tuning—to avoid generating sensitive or non-compliant outputs.
-For example, when asked to explain a recommendation, the model can be instructed not to include health-related inferences or user-specific demographics.
-These filters act as soft constraints that shape the language generation output toward safer and more compliant formats.
-Such behavior can be formalized with a constraint function:
-$P(\text{output contains PII})<\delta$
-where $\delta$ is a defined privacy threshold (ideally close to 0).
-Achieving this objective requires robust prompt design, alignment with safety constraints, and comprehensive evaluation procedures  (Chen et al., 2025a).
-Recent advancements have explored methods such as instruction-driven privacy regularization, rule-guided decoding to suppress sensitive content, and generation pipelines explicitly constrained to avoid exposing personally identifiable information.
-These safeguards are especially critical for large-scale systems that produce user-facing outputs like explanations, summaries, or personalized recommendations in natural language  (Dong et al., 2024b)
-Synthetic Data Generation for Privacy: LLMs can also generate entirely synthetic user behavior data—such as clickstreams, session logs, or user profiles—that mimic the statistical structure of real data without revealing any individual information.
-This synthetic data can be used for training, testing, or debugging recommender systems in compliance with privacy constraints.
-To ensure this approach is effective, synthetic datasets $D_{\text{synth}}$ must satisfy differential privacy guarantees. A typical requirement is:
-$\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\in D]\approx\Pr[\text{LLM outputs }D_{\text{synth}}\mid x\notin D]$
-This ensures that the inclusion or exclusion of any individual data point in the training corpus does not significantly influence the model’s outputs.
-Additional metrics like Jensen-Shannon divergence or Wasserstein distance can be used to compare the distributional similarity between real and synthetic data.
-Recent work has explored differentially private synthetic sequence modeling using transformer-based LLMs (Xie et al., 2024), conditional user profile generation for safe recsys debugging, and adversarially trained privacy-preserving simulators for behavioral logs (Lin et al., 2024).
-These models enable privacy-safe benchmarking and model iteration without relying on production data.
+Balancing innovation with data security requires robust governance frameworks.
+Access control systems must prevent unauthorized use while enabling rapid experimentation.
+Companies such as LinkedIn and Airbnb employ platforms with role-based access control (RBAC), automated policy enforcement, and secure sandboxes (Liu et al., 2021).
+These ensure compliance and accountability without obstructing productivity.
+Data Minimization via LLMs: LLMs can reduce reliance on personally identifiable information (PII) by extracting signals from anonymized data, reviews, or intent prompts, thereby minimizing privacy risks while maintaining personalization:
+[…]
+Privacy-preserving methods include federated prompt learning with differential privacy (Tran et al., 2025), synthetic user simulations, token-level obfuscation, and private fine-tuning (Zhang et al., 2025c).
+Instructable Privacy Filters: LLMs can be guided through prompts or fine-tuning to avoid sensitive outputs, enforcing constraints such as:
+[…]
+where $\delta$ is a privacy threshold. Techniques include privacy regularization, rule-guided decoding, and instruction-driven alignment (Chen et al., 2025a; Dong et al., 2024b).
+Synthetic Data Generation: LLMs can generate synthetic user logs or profiles that mimic real data distributions without exposing individuals. Differential privacy requires:
+[…]
+ensuring individual data points do not affect outputs.
+Metrics like Jensen-Shannon divergence or Wasserstein distance validate distributional similarity.
+Recent work explores DP-based sequence modeling (Xie et al., 2024), conditional profile generation, and adversarial simulators (Lin et al., 2024), enabling privacy-safe benchmarking and debugging.
 5. Limitations & Open Challenges of LLMs in Recommender Systems
@@
 Even with 4-bit weight-only quantization, a 7-B model exceeds 35 ms per query on a single A100 (Rajaraman et al., 2025).
-Token-parallel kernels, KV-cache fusion, and Flash-Attention v3 yield 1.7×1.7\!$\times$1.7 ×2.3×2.3\!$\times$2.3 × speed-ups but degrade under bursts that saturate memory bandwidth (Jo et al., 2025; Shah et al., 2024).
+Token-parallel kernels, KV-cache fusion, and Flash-Attention v3 yield $1.7\!\times$$2.3\!\times$ speed-ups but degrade under bursts that saturate memory bandwidth (Jo et al., 2025; Shah et al., 2024).
 Hierarchical routing—first-stage dense retrieval, second-stage sparse MoE—cuts tail latency by 28 % yet adds routing overhead and fairness concerns (Tairin et al., 2025).
@@
 Workload-aware autoscaling and serverless accelerators lower idle GPU burn but add cold-start penalties that hurt tail latency (Tran et al., 2024).
-Cloud API mark-ups (\$$\sim\$0.03$ per k-tokens) further inflate total cost of ownership in high-throughput pipelines (Rogers et al., 2025).
+Cloud API mark-ups ($\sim\mathdollar 0.03$ per k-tokens) further inflate total cost of ownership in high-throughput pipelines (Rogers et al., 2025).
 Hallucination and Semantic Misalignment – LLMs can fabricate item attributes, user intents, or reasoning chains, surfacing non-existent products or biased analogies (Surana and Thompson, 2023; Patel and Deshmukh, 2024).
@@
 Fidelity critics trained with human feedback reduce error rate another 12 ms latency per call (Shao et al., 2024).
-Negative prompt augmentation and contrastive rejection sampling further reduce hallucination probability to $<4\%$ on internal ad-catalog tests (Liu and Zhu, 2025).
+Negative prompt augmentation and contrastive rejection sampling further reduce hallucination probability to $4\%$ on internal ad-catalog tests (Liu and Zhu, 2025).
 Nonetheless, hallucinations that slip through can propagate via feedback loops, degrading long-term engagement metrics (Pan et al., 2024).
-Prompt Sensitivity and Reproducibility – Microscopic lexical edits (e.g., “recommend” → “suggest”) can produce ranking deltas $>\!10$ NDCG, complicating A/B parity, rollback, and incident triage (Savarese et al., 2020; Mitchell et al., 2019).
+Prompt Sensitivity and Reproducibility – Microscopic lexical edits (e.g., “recommend” → “suggest”) can produce ranking deltas >10>\!10 % NDCG, complicating A/B parity, rollback, and incident triage (Savarese et al., 2020; Mitchell et al., 2019).
 Prompt provenance graphs, canary tokens, and contrastive diff testing improve reproducibility but incur  5 GB/day in lineage logs per million requests (Zhuo et al., 2024).
@@
 Representation Drift and Embedding Incompatibility – Periodic fine-tunes or RLHF updates shift latent spaces, invalidating historical user vectors and cached scores (McNee et al., 2006).
-Embedding versioning, backward-compatible projection heads, and dual-encoder rehearsal maintain cosine similarity $>\!0.85$ across upgrades, but add 2–3 ms per inference and triple index storage (Liu and Rehn, 2025; Qian et al., 2024).
+Embedding versioning, backward-compatible projection heads, and dual-encoder rehearsal maintain cosine similarity >0.85>\!0.85 across upgrades, but add 2–3 ms per inference and triple index storage (Liu and Rehn, 2025; Qian et al., 2024).
 Continual-learning regularizers slow drift but sacrifice 1–2 % relevance on fresh items (Zhao and Meier, 2025).
@@
 Evaluation Gaps – Classic list metrics (NDCG, MAP, Recall@K) ignore semantic fidelity, novelty, and explanation quality central to LLM outputs.
-Hybrid metrics that blend counterfactual click uplift with LLM-judged coherence scores correlate weakly with real satisfaction ($r\!<\!0.35$(Joachims et al., 2018; Yin and Huang, 2025).
+Hybrid metrics that blend counterfactual click uplift with LLM-judged coherence scores correlate weakly with real satisfaction (r<0.35r\!<\!0.35(Joachims et al., 2018; Yin and Huang, 2025).
 LLM-as-judge pipelines reduce annotation cost 10× but inherit biases from the same model family (Bougie and Watanabe, 2025; Wataoka et al., 2024).